### v4

# CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents

Abstract
Semantic code search, retrieving code that matches a given natural language query, is an important task to improve productivity in software engineering. Existing code search datasets face limitations: they rely on human annotators who assess code primarily through semantic understanding rather than functional verification, leading to potential inaccuracies and scalability issues. Additionally, current evaluation metrics often overlook the multi-choice nature of code search. This paper introduces CoSQA+, pairing high-quality queries from CoSQA with multiple suitable codes. We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system. Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 96.4%. Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA. Models trained on CoSQA+ exhibit improved performance. We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
[…]
tcb@breakable
CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents
Jing Gongthanks: Equal Contribution, Yanghui Wu11footnotemark: 1, Linxi Liang11footnotemark: 1, Yanlin Wangthanks: Corresponding Author, Jiachi Chen, Mingwei Liu, Zibin Zheng Sun Yat-sen University {gonj39,wuyh223,lianglx26}@mail2.sysu.edu.cn, yanlin-wang@outlook.com, {chenjch86,liumw26,zhzibin}@mail.sysu.edu.cn,
Introduction
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development  (Li et al., 2013; Sim et al., 2011; Stolee et al., 2014). Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems  (Yan et al., 2020). Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR)  (Husain et al., 2019; Huang et al., 2021; Li et al., 2024a; Lu et al., 2021; Yao et al., 2018; Feng et al., 2020). According to our survey of 102 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 8.31 times per programming day on average. For each query, they typically reference 2.64 code examples to find the necessary information. These findings highlight a key need in creating a multi-choice code search benchmark: effectively pairing real-world queries with multiple correct and relevant code snippets.
Existing code search datasets like CoSQA  (Huang et al., 2021) and CodeSearchNet  (Husain et al., 2019) rely on expert human annotators to evaluate query-code pairs. While this approach captures human judgment, it exists several limitations. First, human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations. Second, manual annotation requires significant time and resources, limiting dataset scalability. Recent works by  Li et al. (2024a) leverage LLMs for annotation to address scalability issues. However, LLMs face challenges in reliably determining whether code functionally fully satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic. Furthermore, most code search benchmarks employ metrics like MRR or Pass@k (Lu et al., 2021), which primarily evaluate single-result retrieval performance. This approach overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations. While alternative metrics like MAP and Recall@k Husain et al. (2019) exist for evaluating multiple code search, they underemphasize the importance of top-ranked results. This is problematic since, according to our survey, users heavily rely on first-page search results.
In this work, we begin with a survey of 102 experienced Python programmers to understand multi-choice code search practices. The survey results reveal that developers frequently require and utilize multiple code examples for a query. Motivated by these insights, we present a new benchmark CoSQA+, with 373,339 pairs of web queries filtered from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search. To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a first screener, test program generator, test executor, bug fixer, and final arbiter. The label indicates whether each code snippet exactly matches its corresponding query requirements. Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks  (Wang et al., 2013). We use it to measure 9 code search methods, 8 of which are embedding-based.
To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 200 randomly selected query-code pairs from CoSQA+. Three Python experts first design test programs for these pairs to establish ground truth annotations. We then compare the accuracies among test-driven agents, python expert annotators without tests, and single LLM. Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 96.4%. Notably, 83.67% of the test programs automatically generated by test-driven agents are executable. Furthermore, using identical data sources (queries from CoSQA and code from CodeSearchNet), we evaluate CoSQA+ against CoSQA as fine-tuning corpora for three models: CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings. Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both NDCG@10 and MRR metrics.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we develop the first Multiagent annotator with over 96.4% accuracy for code search dataset.
  • To the best of our knowledge, we construct CoSQA+ which is the first benchmark designed for evaluating multi-choice code search.
  • We enhance the fully automated construction process through candidate pair construction by multiple models and annotation by agents.
  • We publicly release CoSQA+ to facilitate further research in semantic code search.
Related Work
Code Search Datasets
Besides our work, there are other datasets and benchmarks in the realm of code search.  Husain et al. (2019) present CodeSearchNet which contains a wide range of query-code pairs from GitHub. CoSQA  (Huang et al., 2021) and CS1QA  (Lee et al., 2022) contain the search logs collected from Microsoft Bing search engine but only cover Python and are limited in size. StaQC  (Yao et al., 2018) is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality.  Li et al. (2024b) introduces ProCQA, a large-scale community-based programming question answering dataset mined from StackOverflow with strict filtering strategies for quality and fairness control. This benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios. However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague. Recently,  Li et al. (2024a) present Query4Code, which also leverages LLMs for dataset annotation, but their approach uses a single model for similarity computation and adopts a less strict matching criterion.
Test-Driven Agents
Traditional test generation approaches relied heavily on human expertise, but recent work has demonstrated the potential of automated test agents. LLM-based agents have shown promising results in various testing scenarios. For instance, in the domain of deep learning libraries,  Deng et al. (2024) develops agents capable of generating diverse and meaningful test inputs. Several studies have explored LLM-based agents for specialized tasks. For test oracle generation,  Nashid et al. (2023) utilizes few-shot prompting to create context-aware assertions, achieving a high accuracy. In program repair, studies by  Xia and Zhang (2023) and  Jiang et al. (2023) have demonstrated the effectiveness of integrating LLMs with iterative validation workflows. However, existing test-driven agents primarily focus on validating specific code functionality without addressing the semantic challenges in code search systems, where our test-driven annotation approach bridges this gap through systematic validation of query-code relevance.
Figure 1: The construction process of CoSQA+. The final CoSQA+ dataset includes queries from original CoSQA, and codes filtered from CodeSearchNet.
CoSQA+
Figure 1 illustrates the benchmark construction process for CoSQA+. The process starts with query and code collection, in which we leverage the queries from CoSQA and collect python code snippets filtered from CodeSearchNet. In candidate pairs construction, collected code snippets and queries are embedded as vectors to calculate similarity. For each query, the top 20 query-code pairs with the highest similarity are selected as candidates for potential matches. In model annotation, each query-code pair is annotated by test-driven agents with “1” or “0”, indicating whether the code snippet exactly matches the query. These labeled pairs are then added to CoSQA+.
Query and Code Collection
This section details the reasons why we select CoSQA as the query source and the CodeSearchNet Corpus  (Husain et al., 2019) serves as the code sources.
Query Source
Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes  (Huang et al., 2021). We use the queries as the source for the following reasons: (1) These queries are drawn directly from real-world programming questions, making them representative of actual developer needs and search patterns. For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets  (Sun et al., 2022). (2) Through rigorous filtering and validation processes, the queries demonstrate strong consistency in format and clear programming intent. This careful curation helps maintain high data quality and reduces noise in the training process  (Li et al., 2020). By training on these curated real-world queries, code search models can better align with how developers naturally express their programming needs.
Code Source
To the best of our knowledge, CodeSearchNet is the only origin dataset specifically designed for Python functions in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages  (Husain et al., 2019). According to our survey, code searchers prefer function codes with clear input and output. These codes are high quality with higher readability, testability and maintainability  (Martin, 2009). To extract these high-quality codes from CodeSearchNet’s 1,156,085 Python functions, we filter out 653,994 functions using a rule-based approach. This involves parsing the code into an Abstract Syntax Tree (AST), identifying function nodes with at least one input parameter and a valid return statement. Such testable functions lay the foundation for test-driven agents to design test program and improve dataset quality.
Candidate Query-Code Pairs Construction
Labeling all query-code pairs is impractical. To improve efficiency, we only retain high-confidence pairs before annotation. This section details the construction process of high-confidence query-code pairs.
Construction Steps
The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matching codes from a large codebase. This forms the foundation for annotating query-code pairs in the CoSQA+ dataset. Firstly, we employ three well-established code embedding models to generate vector representations for both queries and codes. Then, these embeddings are used to calculate cosine similarity scores, averaging similarity results across the three models. Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query. For a detailed breakdown of query distributions by the number of exactly matched codes, please see Appendix D. Distribution of Queries Based on Exactly Matched Codes.
Multi-model Embedding
High similarity between query and code is fundamental for a high-quality dataset  (Sun et al., 2022). During our preliminary assessments, we observe disparities in the similarity scores calculated by different models. Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably. And this will potentially lead to unfair test result of benchmark. To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs. Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets. Specifically, we use three models: CodeBERT  (Feng et al., 2020), UniXcoder  (Guo et al., 2022) and CodeT5+ 110M embedding  (Wang et al., 2023).
The relatedness between the query and code is quantified by the cosine similarity, calculated as:
${cosine\_similarity}=\frac{\mathbf{v}_{\text{query}}\cdot\mathbf{v}_{\text{code}}}{\|\mathbf{v}_{\text{query}}\|\|\mathbf{v}_{\text{code}}\|}$
where $\mathbf{v}_{\text{query}}$ and $\mathbf{v}_{\text{code}}$ are the embedding vectors of the query and code, respectively.
By taking the average similarity score from these three models, we promote a more balanced and fair assessment of query-code pair quality.
Top-20 Code Candidates Selection
Before constructing the full set of candidate pairs, we randomly selected 2,000 queries and annotated them using test-driven agents to determine the optimal number of candidates per query. The experimental results indicate that the probability of identifying relevant codes decreases significantly as the number of candidates increases (Figure 2). As shown in our analysis, when the number of candidates reaches 20, the marginal benefit of further expansion diminishes. This decline can be attributed to limitations of codebase, nature of queries and ability of embedding models. Furthermore, adding just one more candidate code per query results in an additional 20,604 pairs, which would incur substantial costs in terms of both time and financial resources for annotation. Thus, limiting candidates to the top-20 strikes a balance between maximizing the retrieval of relevant codes and controlling annotation costs.
Figure 2: Probability of identifying matched codes as the number of candidate codes increases
Test-Driven Agents Annotation
To efficiently and accurately annotate the large number of query-code pairs, we design a test-driven annotation agents based on DeepSeek-V3. An preliminary screener will determine whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment. For uncertain cases, we determine the labels based on the test execution results and final judgment. In both paths, labels are assigned to indicate whether the code matches the query’s requirements (1) or not (0). All labeled query-code pairs are collected to form our dataset CoSQA+, which consists of pairs that have been rigorously verified either through direct assessment or test-based validation. All prompts and a detailed example are provided in AppendixA.
Figure 3: Example illustration of (1) clear match, (2) clear non-match, and (3) unclear cases.
Preliminary Screener
The evaluation process begins with a preliminary screener that categorizes query-code pairs into three distinct labels: (1) clear match, (2) clear non-match, and (3) unclear, as illustrated in Figure 3. Guided by the prompt instructions, the model assesses whether the code fully addresses the query’s functional requirements. For unambiguous cases, it provides a justification and immediate classification. For ambiguous cases (representing 33% of candidates), it flags the pair for deeper verification by downstream agents. This triage mechanism optimizes the annotation workflow by resolving straightforward cases upfront while reserving computational resources for ambiguous scenarios.
Test Program Generator
For ambiguous cases, we employ a test-driven verification methodology. The test program generator constructs executable test suites designed to rigorously validate code functionality. Each test suite includes dependency scaffolding — importing required libraries, defining auxiliary classes, and declaring helper functions — to ensure runtime compatibility. The tests incorporate assert-based validation mechanisms targeting edge cases, functional requirements, and behavioral constraints specified in the query. Considering that a single query may have multiple valid explanations, the test program is generated based on one of the plausible interpretations of the query. This approach transforms ambiguous cases into executable specifications for objective validation.
Test Executor and Bug Fixer
The test executor operates within an isolated Docker environment to ensure safe execution while preventing host system interference. When errors such as ModuleNotFoundError occur, the pipeline initiates a two-stage repair process. First, it attempts direct installation of missing dependencies using standard package managers. If unresolved, the system invokes an LLM to generate context-aware terminal commands for dependency resolution, which are executed iteratively until the error is mitigated or deemed irreparable. After dependency fixes, the test suite is re-executed to validate functionality. Execution traces, assertion outcomes, and error logs are aggregated into a structured report, which is forwarded to the final arbiter for assessment. This approach enhances the reliable verification of ambiguous cases while maintaining environmental integrity, achieving an executable program ratio of 83.67%.
Final Arbiter
The framework ultimately concludes with a final arbiter stage, where the agent employs a specialized prompt to assess if the generated code fully satisfies the query. This assessment is based on both the test program produced and the outcomes of its execution. This methodology shifts verification from theoretical code inspection to empirical observation, empowering the agent to establish evidence-based conclusions by systematically evaluating actual runtime behavior. By grounding judgments in dynamic test execution rather than relying exclusively on static structural analysis and semantic analysis. An example is provided in Appendix BB. An example of Final Arbiter.
Benchmark Statistics
In total, 412,080 labeled query-code pairs are retained. Table 1 shows the statistics of our CoSQA+ benchmark.
Number of pairs Number of code Code average length Average test cases Executable rate
412,080 132,952 378.052 4.31 83.67%
Table 1: Statistics of CoSQA+.
Cost Analysis
For data annotation, a total of 412,080 data samples are labeled by Deepseek-V3 at a cost of approximately \$257, resulting in an average cost of \$0.00062 per sample. The labeling time varies significantly based on complexity: clear cases require approximately 10 seconds per sample, while non-clear cases may take up to 50 seconds, depending on API rate limits, dependency installations, and network latency. However, the annotation process benefits from full concurrency across multiple queries, enabling rapid completion when computational resources are sufficient.
Evaluation
In this section, we conduct experiments with corresponding analysis to answer the following research questions (RQs):
RQ1: Is the quality of CoSQA+ higher than that of CoSQA?
RQ2: Can our test-driven agents surpass Python experts and LLMs in the annotation task?
RQ3: How effective are the existing approaches in multi-choice code search on our benchmark?
RQ1: Quality Comparison between CoSQA+ and CoSQA
We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA.
Metrics
Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG  (Wang et al., 2013) and MRR  (Feng et al., 2020). NDCG measures the quality of ranking by considering both the relevance and position of retrieved results. For a given query $q_{i}$, we calculate NDCG@k as:
$NDCG@k=\frac{DCG@k}{IDCG@k}$
where DCG@k is the Discounted Cumulative Gain up to position k, and IDCG@k is the ideal DCG@k (the maximum possible DCG@k). The final NDCG score is averaged across all queries. MRR evaluates the average reciprocal of the rank at which the first relevant code appears:
$MRR=\frac{1}{|N|}\sum_{i=1}^{|N|}\frac{1}{rank_{i}}$
where |N| is the number of queries and rank_i is the rank of the first relevant code for the i-th query.
Experimental Setup
In the first experiment, we randomly select 1000 query-code pairs labeled "1" from CoSQA+ and CoSQA, ensuring the queries are identical. For CoSQA+, a query may have multiple correct codes, so we randomly select one that differs from the corresponding query’s code in CoSQA. These queries are then presented to experienced programmer volunteers, who are asked to choose the code that satisfies more demand of the query. Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
For the second experiment, we fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with NDCG@10 and MRR as evaluation metrics. From the CSN benchmark we use for validation, known as CSN99, we extract Python test samples and build 399 query-code pairs for 95 queries, referred to as CSN Python. The queries in CSN99 are realistic queries sourced from Bing, carefully selected and verified by humans. In our training process, the model is trained using individual query-code pairs, similar to traditional one-to-one code search models. The CoSQA+ dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division.
Model Metric Fine-tune Dataset
None CoSQA CoSQA+
CodeBERT MRR 0.952 0.939 0.966
NDCG@10 0.958 0.937 0.962
UniXcoder MRR 0.891 0.951 0.954
NDCG@10 0.890 0.946 0.952
CodeT5+ MRR 0.794 0.938 0.940
embedding NDCG@10 0.806 0.943 0.951
Table 2: Performance of code search models fine-tuned with different data. The test set is CSN Python.
Results and Analysis
After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 73.8% of the better codes are from CoSQA+, while 26.2% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table 2, the three models fine-tuned with CoSQA+ outperforms those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in NDCG@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
RQ1 Summary: CoSQA+ has better quality than CoSQA. Fine-tuning CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ yields better performance than with CoSQA, especially for multiple relevant codes.
RQ2: Performance of Agent-based Annotator
We design this experiment to assess whether our agent-based annotation approach can achieve comparable or superior quality to traditional human annotation and single LLM annotation.
Experimental Setup
To establish a reliable ground truth, we first have three Python experts design test programs for 300 randomly selected query-code pairs and annotate them based on test execution results. The final ground truth labels are determined by majority voting among these three experts. We then evaluate three different annotation approaches against this ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet).
For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators.
Results and Analysis
Method Accuracy
Agent-based Annotator 0.964
Human Wo. Test Program 0.890
Deepseek-V3 0.820
O1-mini 0.816
Claude-3.5-Sonnet 0.830
Table 3: Accuracy comparison of different annotation methods against ground truth.
As shown in Table 3, our agent-based annotator achieves 96.4% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (82.0%), O1-mini (81.6%) and Claude-3.5-Sonnet (83.0%). This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification.
Importantly, about 83% of the test programs generated by our agent are successfully executable, indicating the reliability of our test program generation approach. The high executability rate indicates that most query-code pairs can be verified through actual program testing rather than relying solely on semantic analysis.
The superior performance of our agent-based annotators can be attributed to two key factors: 1) The test program provides concrete evidence of whether the code satisfies the query requirements, reducing ambiguity in the annotation process; 2) The systematic verification through test execution helps eliminate human bias and inconsistency that often occur in manual annotation.
In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.85, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
RQ2 Summary: Our agent-based annotator achieves 96.4% accuracy, outperforming both traditional human annotation and single LLM approaches. The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
RQ3: Existing Methods Performance on CoSQA+
This experiment aims to evaluate the effectiveness of different models in accurately retrieving code snippets that match natural language queries on the CoSQA+ benchmark.
Task
The natural language code search task is a text retrieval problem. Given a natural language query $q_{i}$ and a collection of code snippets $C=\{c_{1},\ldots,c_{n}\}$, the goal is to identify all code snippets
$\{c^{*}_{j}\mid c^{*}_{j}\in C\text{ and }c^{*}_{j}\text{ exactly matches }q_{i}\}$
that exactly match the query.
Experimental Setup
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using NDCG@10 as the primary metric. NDCG@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match. Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly.
The tested methods include both traditional information retrieval techniques and advanced deep learning-based models. Specifically, we compare a traditional Bag of Words (BoW) approach (Zhang et al., 2010) with state-of-the-art semantic code search models, namely: CodeBERT (Feng et al., 2020), UniXcoder (Guo et al., 2022), and CodeT5+ embedding (Wang et al., 2023). In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3  (Sturua et al., 2024), All-MiniLM-L12-v2 (Reimers and Gurevych, 2021), All-mpnet-base-v2  (Song et al., 2021), and Multilingual-e5-large  (mul, 2023).
Results and Analysis
Method NDCG@10 MRR
BoW 0.046 0.034
CodeBERT 0.232 0.175
UniXcoder 0.187 0.140
CodeT5+ embedding 0.179 0.136
Jina-embedding-v3 0.225 0.171
All-MiniLM-L12-v2 0.214 0.163
All-mpnet-base-v2 0.227 0.172
Multilingual-e5-large 0.218 0.166
Table 4: Performance of Various Methods on CoSQA+ Using NDCG@10 and MRR Metrics.
While traditional information retrieval methods, such as BoW, are computationally efficient, Table 4 demonstrates that they struggle to capture the semantic subtleties needed to accurately map natural language queries to relevant code snippets, particularly when multiple valid solutions exist. Among the deep learning-based models, CodeBERT achieves the highest performance with an NDCG@10 of 0.232 and an MRR of 0.175. This strong performance can be attributed to its encoder-only architecture and extensive pre-training on large-scale paired natural and programming language corpora, especially Python, which enhances its code understanding capabilities. Other embedding-based models also show substantial improvements over traditional methods, highlighting the benefits of modern deep learning architectures for semantic code search.
RQ3 Summary: The evaluation on CoSQA+ clearly indicates that deep learning-based models significantly outperform traditional retrieval methods for semantic code search. Notably, CodeBERT leads the performance ranking by achieving the highest scores in both NDCG@10 and MRR, while the alternative embedding-based models also offer competitive results.
Conclusion
In this paper, we present CoSQA+, the first large-scale multi-choice code search dataset with 412,080 pairs. Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks. Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA. The test-driven annotation approach achieves a high accuracy of 96.4%, outperforming traditional human annotation and single LLM methods. Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks. We believe our agents-annotated CoSQA+ benchmark will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task
Ethical consideration
In this work, we prioritize ethical considerations by the responsible use of data and methodologies. The dataset sources, CoSQA and CodeSearchNet, are publicly available under permissible licenses, mitigating privacy concerns. Our test-driven agents are designed to minimize human bias by automating code evaluation, thus enhancing fairness and consistency. Additionally, we openly release CoSQA+ to promote transparency and reproducibility in code search research. While our approach aims to improve code search efficiency, we acknowledge the potential for misuse in generating or retrieving malicious code. We encourage the community to adopt ethical guidelines when using our benchmark and tools, ensuring that their application aligns with positive social impact.
Limitations
We are committed to using LLM-based test-driven annotation to construct high-quality function-level multi-choice code search datasets, but there are still several limitations. Firstly, due to resource constraints, we only apply the construction method to Python code. The method is generalizable, and we plan to extend it to other programming languages. Secondly, we conduct extensive experiments, research, and manual validation to demonstrate the effectiveness of the construction method and the quality of the dataset, but due to limitations in time and resources, CoSQA+ may contain some biases or ignore some scenarios. Finally, although various mechanisms are designed to improve the success rate and accuracy of automated annotation and achieve promising results, factors such as ambiguity in query statements, complexity of code functionality, and uncertainty in program execution environments prevent ideal outcomes for few query-code pair annotations. We plan to continue exploring and improving the framework and methods for automated annotation in the future.
Appendix A A. Prompts for Test-Driven Agents
Preliminary Screener
1def get_preliminary_screening_prompt( query, function_code):
2 prompt = f"""
3 Instructions for Evaluation:
4 You are provided with a query that specifies requirements for a function and a block of function code.
5 Your goal is to determine if the code fully matches the query’s requirements.
6 - If the code clearly does not satisfy the query, output preliminary_screening: 0 and provide a corresponding explanation.
7 - If the code clearly satisfies the query, output preliminary_screening: 1 and provide a corresponding explanation.
8 - If the determination is ambiguous or if further testing using a test_program_generation agent is needed, output preliminary_screening: 0.5 and provide a corresponding explanation.
9
10 Ensure your answer is formatted exactly as follows:
11 preliminary_screening: [0/0.5/1],
12 explanation: <Your explanation in English>
13
14
[…]
17 """
18 return prompt
Test Program Generator
1def get_test_program_prompt(query, function_code):
2 prompt = f"""
3 Task Description:
4 Please generate a test program based on the provided query and function code. The test program should:
5 1. Contain assert statements in the main function to verify that the function code correctly implements the behavior described in the query.
6 2. **Do not** test other classes or functions generated, only focus on testing the behavior specified in the query.
7 3. The generated test program should be complete, executable, and must test the functionality described by the query using assert statements.
8
9 For example, given the following query and function code:
10
11 query: "python remove all empty items in list"
12 code:
13 def remove_from_list(list, remove_items):
14 list = return_list(list)
15 remove_items = return_list(remove_items)
16 for remove in remove_items:
17 if remove in list:
18 list.remove(remove)
19 return list
20
21 The generated program should:
22 Test the behavior described in the query (removing empty items from a list).
23 Include only the assert statements in the main function.
24 Be a fully executable test program that can verify the behavior of function code.
25
26 Example Test Program Output:
27
28 # Additional function definition
29 def return_list(input_list):
30 return [item for item in input_list if item]
31
32 # Original function code
33 def remove_from_list(list, remove_items):
34 list = return_list(list)
35 remove_items = return_list(remove_items)
36 for remove in remove_items:
37 if remove in list:
38 list.remove(remove)
39 return list
40
41 # Test code
42 def main():
43 # Test case 1: Remove empty items from the list
44 list1 = [1, 2, "", 4, None, "test", ""]
45 remove_items1 = [""]
46 result1 = remove_from_list(list1, remove_items1)
47 # Assert that empty strings have been removed from the list
48 assert result1 == [1, 2, 4, None, "test"]
49
50 # Test case 2: Remove None from the list
51 list2 = [1, 2, "", 4, None, "test", ""]
52 remove_items2 = [None]
53 result2 = remove_from_list(list2, remove_items2)
54 # Assert that None has been removed from the list
55 assert result2 == [1, 2, "", 4, "test", ""]
56
57 # Test case 3: Ensure no changes when removing an item not in the list
58 list3 = [1, 2, 3]
59 remove_items3 = [4]
60 result3 = remove_from_list(list3, remove_items3)
61 # Assert that the list remains unchanged
62 assert result3 == [1, 2, 3]
63
64 # Test case 4: Empty list case
65 list4 = []
66 remove_items4 = [""]
67 result4 = remove_from_list(list4, remove_items4)
68 # Assert that the result is still an empty list
69 assert result4 == []
70
71 print("All tests passed.")
72
73 if __name__ == "__main__":
74 main()
75
76 Now please directly and only generate the complete test program based on the instructions above. Do NOT add any explanation.
77 query: {query}
78 code: {function_code}
79 """
80
81 return prompt
Bug Fixer
1def get_pip_install_command_prompt( test_output: str, test_program: str) -> str:
2 prompt=f"""You are a Python development assistant. The user encountered a ModuleNotFoundError when running their test code. Here are the details:
3 Test Output:
4 {test_output}
5
6 Test Program:
7 {test_program}
8
9 Please determine the correct pip install command(s) required to install the missing module(s). Consider that the PyPI package name might differ from the module name. Provide only the necessary pip install command(s), each on a separate line. Do not include any other text or explanation.
10
11 Examples of valid responses:
12 1. If missing ’cv2’:
13 pip install opencv-python
14
15 2. If missing ’yaml’:
16 pip install pyyaml
17
18 3. If multiple packages needed:
19 pip install numpy
20 pip install pandas
21
22 Now, provide the correct pip install command(s) for the current error:
23 """
24 return prompt
Final Arbiter
1 def get_final_judge_prompt( function_code: str, query: str, test_program: str, test_result: str,
2 test_output: str) -> str:
3 prompt = f"""
4 You are an code expert in judging whether a function code can fully satisfy the requirement of the query. You can refer to a generated test program and its test result and test output.
5
6 You are provided with the following inputs:
7
8 1. Query:
9 {query}
10
11 2. Function Code:
12 {function_code}
13
14 3. Test Program:
15 {test_program}
16
17 4. Test Result:
18 {test_result}
19
20 5. Test Output:
21 {test_output}
22
23 Your Task:
24 Evaluate whether the provided function code completely satisfies the query’s requirements based on a detailed analysis of the function code and the results and the output from running the test program.
25
26 Judgment Guidelines:
27 - If the function code clearly fails to meet the query requirements (supported by incorrect logic or failing test results), then output:
28 final_verdict: 0,
29 final_explanation: "A clear explanation describing why the function code does not meet the query requirements, referencing any issues in the code and test failures."
30
31 - If the function code fully meets all the query requirements (supported by correct logic and successful test results), then output:
32 final_verdict: 1,
33 final_explanation: "A clear explanation describing how the function code meets all the query requirements, referencing the successful test outcomes and correct implementation."
34
35 Important:
36 - Your final output must strictly adhere to the following format:
37 final_verdict: [0 or 1],
38 final_explanation: <Your explanation in English>
39 - Only output 0 or 1 as the final verdict.
40
41 Please provide your final judgment as instructed.
42 """
43 return prompt
B. An example of Final Arbiter
This figure illustrates the process of handling an unclear case. The test program generator generates a program, which is then executed by the test executor. The test results are returned to the final arbiter, which assigns a label to the case.
Figure 4: Process of Handling an Unclear Case through Test Program Generation, Execution, and Final Arbitration
C. Code Search Behavior Survey
We conduct this survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.
Survey Demographics and Methodology
A total of 102 participants complete the survey, predominantly undergraduate students (89.22%) with some master’s students (8.82%). The majority of participants (92.16%) have formal computer science education, and most have 1-5 years of programming experience (66.67%), while 30.39% have less than one year of experience. All participants are Chinese university students who use Python for development, with 52.94% considering Python as their primary programming language. The survey is conducted anonymously through an online platform, with all questions presented in Chinese. No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
Survey Results and Analysis
Our survey reveal several significant patterns in code search behavior. Most notably, when 66.67% typically needing 2-3 code examples to understand and implement a solution. Participants evaluate search results based on several key factors, primarily focusing on implementation relevance (65.69%) and solution differences (56.86%). When presented with multiple code examples, they particularly value the clarity of implementation differences and the completeness of edge case handling.
Survey Questions
Table 5 shows the key questions from our survey (translated from Chinese).
Questions Response Rate (%)
A B C D
How often do you perform code searches during programming-related tasks per day?
(A: >10 times, B: 6-10, C: 3-5, D: 1-2) 50.00 19.61 14.71 15.68
In your daily development, how many code examples do you typically need to reference for a query (such as API usage, algorithm implementation)
(A: Only 1, B: 2-3, C: 4-5, D: >5) 16.67 66.67 12.75 3.92
In which scenarios do you most frequently need to reference multiple code examples when performing code searches? [Multiple Choice]
(A: API usage, B: Algorithm implementation, C: Performance optimization solutions, D: Other) 68.63 60.78 51.96 47.06
When performing code searches, which better matches your habits?
(A: Only need to find one code snippet that meets the requirements, B: Need to find more than one snippet that meets the requirements) 49.02 50.98 - -
When search returns multiple code examples, which aspects do you focus on most? [Multiple Choice]
(A: Differences between code examples, e.g. comparison between different implementation methods, B: Relevance of each example, e.g. degree of matching with the query, C: Whether the display order of examples meets expectations, e.g. most relevant first, D: Whether examples cover common use cases or edge conditions) 56.86 65.69 49.02 38.24
When evaluating code search results, which do you prioritize more?
(A: Position of the first result that meets requirements, B: Average position of the first few results that meet requirements, C: Diversity of results, D: Other) 44.12 37.25 18.63 0.00
When searching for code examples, which type of function structure do you prefer?
(A: Functions with both input and output, B: Functions with input but no output, C: Functions with output but no input , D: Functions with neither input nor output) 71.57 16.67 8.82 2.94
When searching for code examples, which type of function structure do you end up using?
(A: Functions with both input and output, B: Functions with input but no output, C: Functions with output but no input , D: Functions with neither input nor output) 62.75 22.55 11.76 2.94
Table 5: Distribution of survey responses (n=102). For each question, the options are listed in parentheses.
D. Distribution of Queries Based on Exactly Matched Codes
The table 6 illustrates the distribution of queries according to the number of exactly matched codes, providing insight into the frequency of queries with varying code match counts.
Table 6: Distribution of Queries Based on Exactly Matched Codes
Number of Exactly Matched Codes (n) Query Count
1 2907
2 2445
3 2121
4 1748
5 1429
6 1266
7 1108
8 873
9 724
10 628
11 481
12 398
13 345
14 220
15 155
16 126
17 78
18 44
19 19
20 3

### v5

@@
-CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents
+CoSQA+: Enhancing Code Search Evaluation with a Multi-Choice Benchmark and Test-Driven Agents
 Abstract
@@
 We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system.
-Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 96.4%.
+Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 92.0%.
 Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA.
@@
 We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
-[…]
-tcb@breakable
-CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents
-Jing Gongthanks: Equal Contribution, Yanghui Wu11footnotemark: 1, Linxi Liang11footnotemark: 1, Yanlin Wangthanks: Corresponding Author, Jiachi Chen, Mingwei Liu, Zibin Zheng Sun Yat-sen University {gonj39,wuyh223,lianglx26}@mail2.sysu.edu.cn, yanlin-wang@outlook.com, {chenjch86,liumw26,zhzibin}@mail.sysu.edu.cn,
-Introduction
-Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development  (Li et al., 2013; Sim et al., 2011; Stolee et al., 2014).
-Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems  (Yan et al., 2020).
+Index Terms:
+software engineering, information search and retrieval, human-computer interaction.
+11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
+I Introduction
+Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development [1, 2, 3, 4, 5].
+Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems [6].
 Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories.
-Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR)  (Husain et al., 2019; Huang et al., 2021; Li et al., 2024a; Lu et al., 2021; Yao et al., 2018; Feng et al., 2020).
+Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR) [7, 8, 9, 10, 11, 12].
 According to our survey of 102 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 8.31 times per programming day on average.
@@
 These findings highlight a key need in creating a multi-choice code search benchmark: effectively pairing real-world queries with multiple correct and relevant code snippets.
-Existing code search datasets like CoSQA  (Huang et al., 2021) and CodeSearchNet  (Husain et al., 2019) rely on expert human annotators to evaluate query-code pairs.
-While this approach captures human judgment, it exists several limitations.
-First, human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations.
-Second, manual annotation requires significant time and resources, limiting dataset scalability.
-Recent works by  Li et al.
-(2024a) leverage LLMs for annotation to address scalability issues.
-However, LLMs face challenges in reliably determining whether code functionally fully satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic.
-Furthermore, most code search benchmarks employ metrics like MRR or Pass@k (Lu et al., 2021), which primarily evaluate single-result retrieval performance.
-This approach overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations.
-While alternative metrics like MAP and Recall@k Husain et al.
-(2019) exist for evaluating multiple code search, they underemphasize the importance of top-ranked results.
-This is problematic since, according to our survey, users heavily rely on first-page search results.
-In this work, we begin with a survey of 102 experienced Python programmers to understand multi-choice code search practices.
-The survey results reveal that developers frequently require and utilize multiple code examples for a query.
-Motivated by these insights, we present a new benchmark CoSQA+, with 373,339 pairs of web queries filtered from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search.
-To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a first screener, test program generator, test executor, bug fixer, and final arbiter.
+However, existing approaches face significant challenges in scalability, annotation quality, and evaluation metrics:
+  1. C1
+    Scalability Issues in Human-Based Benchmarking.
+Conventional benchmarks like CoSQA [8] and CodeSearchNet (CSN) [7] rely on expert human annotators to evaluate query-code pairs.
+While this approach captures human judgment, manual annotation requires significant time and resources, limiting dataset scalability.
+The labor-intensive nature of this process makes it impractical for creating large-scale benchmarks needed for robust code search evaluation.
+  2. C2
+    Accuracy Limitations in Annotation Methods.
+Both human and LLM-based annotation approaches suffer from accuracy issues.
+Human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations.
+While Li et al.
+leverage LLMs to address scalability issues [9], these models struggle to reliably determine whether code functionally satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic.
+This fundamental limitation affects the quality of both human and LLM-based annotations [13].
+  3. C3
+    Inadequate Evaluation Metrics for Code Search.
+Most code search benchmarks employ metrics like MRR or Pass@k [10] which primarily evaluate single-result retrieval performance.
+This overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations.
+While alternative metrics like MAP and Recall@k [7] exist for evaluating multiple code matches, they underemphasize the importance of top-ranked results.
+This is particularly problematic as our survey reveals users heavily rely on first-page search results, highlighting the need for better evaluation metrics.
+To address these limitations, we present a new benchmark CoSQA+, with 412,080 pairs of web queries from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search.
+To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a preliminary screener, a test program generator, a test executor, a bug fixer, and a final arbiter.
 The label indicates whether each code snippet exactly matches its corresponding query requirements.
-Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks  (Wang et al., 2013).
-We use it to measure 9 code search methods, 8 of which are embedding-based.
-To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 200 randomly selected query-code pairs from CoSQA+.
+Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks [14].
+We use it to measure 8 code search methods, 7 of which are embedding-based.
+To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 400 randomly selected query-code pairs from CoSQA+.
 Three Python experts first design test programs for these pairs to establish ground truth annotations.
-We then compare the accuracies among test-driven agents, python expert annotators without tests, and single LLM.
-Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 96.4%.
+We then compare the accuracies among test-driven agents, python expert annotators without tests, and a single LLM.
+Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 92.0%.
 Notably, 83.67% of the test programs automatically generated by test-driven agents are executable.
@@
 Our contributions can be summarized as follows:
-  • To the best of our knowledge, we develop the first Multiagent annotator with over 96.4% accuracy for code search dataset.
+  • To the best of our knowledge, we develop the first Multiagent annotator with 92.0% accuracy for code search dataset.
   • To the best of our knowledge, we construct CoSQA+ which is the first benchmark designed for evaluating multi-choice code search.
@@
   • We publicly release CoSQA+ to facilitate further research in semantic code search.
-Related Work
-Code Search Datasets
-Besides our work, there are other datasets and benchmarks in the realm of code search.
-Husain et al.
-(2019) present CodeSearchNet which contains a wide range of query-code pairs from GitHub.
-CoSQA  (Huang et al., 2021) and CS1QA  (Lee et al., 2022) contain the search logs collected from Microsoft Bing search engine but only cover Python and are limited in size.
-StaQC  (Yao et al., 2018) is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality.
-Li et al.
-(2024b) introduces ProCQA, a large-scale community-based programming question answering dataset mined from StackOverflow with strict filtering strategies for quality and fairness control.
+II Motivation Survey
+Do developers really need multi-choice code search?
+To answer this question, we conduct a survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.Fig.
+1 provides a multi-choice code search example
+Figure 1: An example of multi-choice code search. The query is vague in its definition of "illegal character," allowing eight candidate code snippets to match under different interpretations.
+Methodology
+The survey targets 102 participants, predominantly undergraduate students (89.22%) with a smaller proportion of master’s students (8.82%).
+Most of the participants (92.16%) have formal computer science education, and their experience in Python programming varies.
+The majority (66.67%) have between 1 and 5 years of programming experience, while 30.39% have less than one year of experience.
+The participants are selected from programmers using Python as their primary programming language (52.94%).
+The survey is conducted anonymously through an online platform.
+No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
+Results and Analysis
+Our survey reveals several significant patterns in code search behavior.
+Most notably, when 66.67% typically needing 2-3 code examples to understand and implement a solution.
+This suggests that the current code search task, which typically prioritizes a single result, may need to evolve to provide multiple relevant code snippets.
+Participants evaluate search results based on several key factors, primarily focusing on implementation relevance (65.69%) and solution differences (56.86%).
+These preferences emphasize that developers do not just want the closest match to their query, but also value diverse approaches that help them understand the problem and solution better.
+When presented with multiple code examples, they particularly value the clarity of implementation differences and the completeness of edge case handling.
+III Related Work
+III-A Code Search Datasets
+Besides our work, there are other text-to-code datasets and benchmarks in the realm of code search.
+They can be categorized into three types based on the origin of the queries: doc-code datasets, question-code datasets, and query-code datasets.
+CodeSearchNet [7], CoDesc [15], PyMT5 [16], and XLCoST [17] are doc-code datasets, which contain massive code snippets from public repositories with corresponding comments.
+They lay foundational training grounds for code search models to grasp basic code-comment relationships.
+CodeSearchNet [7] dataset comprises a large-scale collection of query-code pairs extracted from GitHub repositories.
+CoDesc [15] is a comprehensive dataset featuring 4.2 million Java methods paired with natural language descriptions, meticulously filtered to minimize noise.
+PyMT5, on the other hand, offers a vast collection of 26 million Python methods and 7.7 million method-docstring pairs, enabling bidirectional translation between code and docstrings.
+XLCoST [17] stands out as the largest multilingual parallel dataset, encompassing 8 languages.
+It supports 10 cross-lingual code tasks, such as translation and search, and provides essential baselines for model development in diverse programming contexts.
+StaQC [11], CoNaLa [18], SO-DS [19], CS1QA [20], and ProCQA [21] are question-code datasets, which contain code answers with corresponding questions from Stack Overflow.
+They challenge models to excel in practical problem-solving scenarios by understanding complex developer inquiries.
+StaQC [11] is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality.
+CoNaLa [18] is a curated dataset derived from Stack Overflow, comprising 2,379 training and 500 test examples of natural language-code pairs, alongside a large automatically-mined dataset of 600k examples, designed for tasks like code synthesis, retrieval, and summarization.
+SO-DS [19], a high-quality snippet collection mined from highly upvoted Stack Overflow posts, focuses on annotated code search, offering 12,137 snippets and 7,674 unique descriptions, with 2,225 annotated queries for validation and testing, specifically tailored for improving code retrieval systems by leveraging natural language descriptions.
+CS1QA [20], a dataset for code-based question answering in programming education, comprises 9,237 annotated question-answer pairs from Python chat logs, alongside 17,698 unannotated examples, designed to benchmark models on code comprehension and question answering in an educational context.
+ProCQA [21] introduces a large-scale community-based programming question answering dataset, which is mined from StackOverflow with strict filtering strategies for quality and fairness control.
 This benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios.
 However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague.
-Recently,  Li et al.
-(2024a) present Query4Code, which also leverages LLMs for dataset annotation, but their approach uses a single model for similarity computation and adopts a less strict matching criterion.
-Test-Driven Agents
-Traditional test generation approaches relied heavily on human expertise, but recent work has demonstrated the potential of automated test agents.
+CoSQA [8], CSN99 of CodeSearchNet [7], WebQueryTest of CodeXGLUE [10], and Query4Code [9] are query-code datasets, which contain real human queries from web and code snippets matched with them.
+They are crucial for training and testing models’ capabilities in handling authentic, diverse search queries.
+CoSQA [8] introduces a dataset of 20,604 human-annotated natural language query-code pairs, enhancing semantic matching for code search, and proposes CoCLR, a contrastive learning method that boosts code question answering accuracy by 10.5% when combined with CodeBERT.
+CSN99, part of the CodeSearchNet Challenge [7], offers 99 natural language queries with approximately 4k expert-annotated relevance judgments, facilitating robust evaluation of code search models.
+The queries of CoSQA and CSN99 are collected from Microsoft Bing search engine.
+WebQueryTest [10] addresses the gap between real user search queries and existing code search datasets by providing a testing set of about 1k real web queries paired with Python code functions, classified for code-search intent and query-code relevance, with annotations from 13 proficient developers.
+Query4Code leverages Large Language Models (LLMs) to annotate datasets, primarily relying on the semantic information embedded within the code [9].
+However, it lacks test cases or validation mechanisms to ensure the accuracy and reliability of the annotations.
+Based on these datasets, COIR [22] emerges as a robust and comprehensive benchmark specifically designed for code retrieval.
+COIR comprises ten curated datasets spanning eight retrieval tasks across seven diverse domains, offering a versatile evaluation platform for state-of-the-art retrieval models.
+Its user-friendly Python framework, compatible with popular benchmarks like MTEB [23] and BEIR [24], facilitates seamless integration into research workflows.
+Despite the advancements of these code search datasets, several limitations persist.
+Doc-code datasets oversimplify the search task by focusing on direct comment-to-code mappings, lacking the complexity of real-world queries.
+Question-code datasets are not designed for code search tasks, emphasizing question-answer context over direct code retrieval.
+Query-code datasets, though highly relevant, are constrained by their difficulty in constructing and annotating at scale.
+III-B Code Search Methods
+The existing methods for code search can mainly be divided into two mainstreams: information retrieval (IR) based methods and deep learning (DL) based models.
+IR-based methods extract keywords from queries and search for these keywords in the code base to retrieve the most relevant code snippets [25, 26, 27].
+However, standard information retrieval methods have drawbacks in the code search domain, as they treat codes as text while programming languages and natural languages differ greatly.
+Fortunately, DL-based methods receive great attention in NLP and achieve significant breakthroughs in code search research.
+The CodeSearchNet [7] dataset paves the way for CodeBERT [12] and GraphCodeBERT [28], both of which are bimodal pre-trained BERT [29] based models that leverage Masked Language Modeling (MLM) and Replaced Token Detection (RTD) [30] objective.
+UniXcoder [31] unifies three pre-training designs into one architecture and utilizes AST structure and code comment to enhance the cross-modal alignment.
+Recent T5 [32] based models like CodeT5 [33] and CodeT5+ [34] show outstanding performance in code search, utilizing the text-to-text transformer architecture and pre-training on large code datasets.
+While the two models achieve high performances in the CSN benchmark, GraphCodeBERT [28] demonstrates superior results by harnessing data flow encoded in the Abstract Syntax Tree (AST) of code to enrich the structural information during pre-training.
+Recent advancements in text embedding models have significantly enhanced the performance of code search methods, particularly in multilingual and long-context retrieval tasks.
+Among these, jina-embeddings-v3 [35] stands out as a state-of-the-art model with 570 million parameters, achieving superior performance on the MTEB benchmark.
+Its innovative use of task-specific Low-Rank Adaptation (LoRA) [36] adapters enables high-quality embeddings for various tasks, including query-document retrieval and text matching, while supporting context lengths of up to 8192 tokens.
+This model outperforms both proprietary embeddings like OpenAI and Cohere.
+multilingual-e5-large [37, 38], which is trained on 1 billion multilingual text pairs and fine-tuned on labeled datasets, is efficient, especially in reducing embedding dimensions without compromising quality, thanks to Matryoshka Representation Learning [39].
+But it does not match the flexibility and performance of jina-embeddings-v3.
+On the other hand, all-MiniLM-L12-v2 [40] and all-mpnet-base-v2 [41] represent efficient, compressed models derived from larger pre-trained Transformers.
+The all-MiniLM-L12-v2 model, based on MiniLM [42], employs deep self-attention distillation to retain high accuracy with reduced parameters, making it suitable for latency-sensitive applications.
+Similarly, all-mpnet-base-v2, built on MPNet [43], combines the strengths of BERT and XLNet [44] by leveraging masked and permuted language modeling, achieving competitive results on downstream tasks [45].
+While these models are effective for general-purpose embeddings, they lack the multilingual and long-context capabilities of jina-embeddings-v3 and multilingual-e5-large, limiting their applicability in complex code search scenarios.
+III-C Test-Driven Agents
+Traditional test generation approaches rely heavily on human expertise, but recent work demonstrates the potential of automated test agents.
 LLM-based agents have shown promising results in various testing scenarios.
-For instance, in the domain of deep learning libraries,  Deng et al.
-(2024) develops agents capable of generating diverse and meaningful test inputs.
-Several studies have explored LLM-based agents for specialized tasks.
-For test oracle generation,  Nashid et al.
-(2023) utilizes few-shot prompting to create context-aware assertions, achieving a high accuracy.
-In program repair, studies by  Xia and Zhang (2023) and  Jiang et al.
-(2023) have demonstrated the effectiveness of integrating LLMs with iterative validation workflows.
-However, existing test-driven agents primarily focus on validating specific code functionality without addressing the semantic challenges in code search systems, where our test-driven annotation approach bridges this gap through systematic validation of query-code relevance.
-Figure 1: The construction process of CoSQA+. The final CoSQA+ dataset includes queries from original CoSQA, and codes filtered from CodeSearchNet.
-CoSQA+
-Figure 1 illustrates the benchmark construction process for CoSQA+.
+For instance, in the domain of deep learning libraries, agents capable of generating diverse and meaningful test inputs are proposed [46, 47], alongside other studies exploring LLM-based agents for specialized tasks.
+For test oracle generation, few-shot prompting is used to generate context-aware assertions for test oracle generation [48, 49], achieving a high accuracy.
+In program repair, some studies [50, 51, 52, 53, 54] demonstrate the effectiveness of integrating LLMs with iterative validation workflows [55].
+However, existing test-driven agents primarily focus on validating specific code functionality without addressing the semantic challenges in code search systems, where our test-driven annotation approach bridges this gap through systematic validation of query-code relevance [56].
+Figure 2: The construction process of CoSQA+.
+(a) Queries sourced from CoSQA and code snippets filtered from CodeSearchNet.
+(b) Candidate pairs selected based on query-code cosine similarity, computed using a multi-model approach.
+(c) Pipeline for annotation by test-driven agents.
+IV CoSQA+
+Fig.
+2 illustrates the benchmark construction process for CoSQA+.
 The process starts with query and code collection, in which we leverage the queries from CoSQA and collect python code snippets filtered from CodeSearchNet.
@@
 These labeled pairs are then added to CoSQA+.
-Query and Code Collection
-This section details the reasons why we select CoSQA as the query source and the CodeSearchNet Corpus  (Husain et al., 2019) serves as the code sources.
+IV-A Query and Code Collection
+This section details the reasons why we select CoSQA as the query source and the CodeSearchNet Corpus [7] serves as the code sources.
 Query Source
-Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes  (Huang et al., 2021).
+Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes [8].
 We use the queries as the source for the following reasons: (1) These queries are drawn directly from real-world programming questions, making them representative of actual developer needs and search patterns.
-For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets  (Sun et al., 2022).
+For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets [57].
 (2) Through rigorous filtering and validation processes, the queries demonstrate strong consistency in format and clear programming intent.
-This careful curation helps maintain high data quality and reduces noise in the training process  (Li et al., 2020).
+This careful curation helps maintain high data quality and reduces noise in the training process [58].
 By training on these curated real-world queries, code search models can better align with how developers naturally express their programming needs.
@@
 While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet.
-The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages  (Husain et al., 2019).
+The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages [7].
 According to our survey, code searchers prefer function codes with clear input and output.
-These codes are high quality with higher readability, testability and maintainability  (Martin, 2009).
+These codes are high quality with higher readability, testability and maintainability [59].
 To extract these high-quality codes from CodeSearchNet’s 1,156,085 Python functions, we filter out 653,994 functions using a rule-based approach.
@@
 Such testable functions lay the foundation for test-driven agents to design test program and improve dataset quality.
-Candidate Query-Code Pairs Construction
+IV-B Candidate Pairs Construction
 Labeling all query-code pairs is impractical.
@@
 Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query.
-For a detailed breakdown of query distributions by the number of exactly matched codes, please see Appendix D.
-Distribution of Queries Based on Exactly Matched Codes.
+Fig.
+3 shows a detailed breakdown of query distributions by the number of exactly matched codes.
+Figure 3: Query counts grouped by number of matched codes
 Multi-model Embedding
-High similarity between query and code is fundamental for a high-quality dataset  (Sun et al., 2022).
+High similarity between query and code is fundamental for a high-quality dataset [57].
 During our preliminary assessments, we observe disparities in the similarity scores calculated by different models.
 Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably.
-And this will potentially lead to unfair test result of benchmark.
+This potentially leads to unfair test results of the benchmark.
 To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs.
+We empirically validate this practice.
+Specifically, we use a single model to select 20 candidate codes for each query and use an ensemble of models, CodeBERT [12], UniXcoder [31] and CodeT5+ 110M embedding [34], to select another 20 candidates.
 Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets.
-Specifically, we use three models: CodeBERT  (Feng et al., 2020), UniXcoder  (Guo et al., 2022) and CodeT5+ 110M embedding  (Wang et al., 2023).
+We calculate the overlap between the candidates chosen by the single model and those chosen by the ensemble and analyze the overlap between them.
+Although bias cannot be completely eliminated, the experiment result in Table I shows that comprehensively utilizing multiple models for matching can effectively reduce the bias of query-code pairs toward any single model.
+As can be seen, the proportion of overlapping codes is low, indicating that the codes matched by multiple models for a query do not show a significant preference for any particular single model.
+TABLE I: Overlap ratio of candidate codes between individual models and the ensemble of models.
+Model Overlap Ratio
+CodeBERT 0.135
+UniXcoder 0.164
+CodeT5+ embedding 0.088
 The relatedness between the query and code is quantified by the cosine similarity, calculated as:
@@
 where $\mathbf{v}_{\text{query}}$ and $\mathbf{v}_{\text{code}}$ are the embedding vectors of the query and code, respectively.
-By taking the average similarity score from these three models, we promote a more balanced and fair assessment of query-code pair quality.
+By taking the average similarity score from these three models, we achieve a more balanced and fair assessment of query-code pair quality.
 Top-20 Code Candidates Selection
-Before constructing the full set of candidate pairs, we randomly selected 2,000 queries and annotated them using test-driven agents to determine the optimal number of candidates per query.
-The experimental results indicate that the probability of identifying relevant codes decreases significantly as the number of candidates increases (Figure 2).
+Before constructing the full set of candidate pairs, we randomly select 2,000 queries and annotated them using test-driven agents to determine the optimal number of candidates per query.
+The experimental results indicate that the probability of identifying relevant codes decreases significantly as the number of candidates increases (Fig.
+4).
 As shown in our analysis, when the number of candidates reaches 20, the marginal benefit of further expansion diminishes.
@@
 Figure 4: Probability of identifying matched codes as the number of candidate codes increases
-Test-Driven Agents Annotation
-To efficiently and accurately annotate the large number of query-code pairs, we design a test-driven annotation agents based on DeepSeek-V3.
-An preliminary screener will determine whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment.
+IV-C Test-Driven Agents Annotation
+To efficiently and accurately annotate the large number of query-code pairs, we design the test-driven annotation agents based on DeepSeek-V3 [60], using the algorithm shown in Algorithm 1.
+Initially, A preliminary screener determines whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment.
 For uncertain cases, we determine the labels based on the test execution results and final judgment.
@@
 All labeled query-code pairs are collected to form our dataset CoSQA+, which consists of pairs that have been rigorously verified either through direct assessment or test-based validation.
-All prompts and a detailed example are provided in AppendixA.
+Algorithm 1 Test-Driven Agent Annotation
+Input: query $q$, code snippet $c$, max_try_time
+Output: label $l\in\{0,1\}$
+1:  $assessment\leftarrow\text{preliminary\_screener}(q,c)$
+[…]
+3:     return  1
+[…]
+5:     return  0
+[…]
+7:     $test\_prog\leftarrow\text{generate\_test\_program}(q,c)$
+8:     $exec\_res\leftarrow\text{execute\_test}(test\_prog)$
+9:     $try\_time\leftarrow 0$
+10:     while $exec\_res.\text{error}\ \textbf{and}{}\ try\_time\leq max\_try\_time$ do
+11:        $terminal\_commands\leftarrow\text{fix\_test\_program}(test\_prog,exec\_res.\text{error})$
+[…]
+13:        $exec\_res\leftarrow\text{execute\_test}(test\_prog)$
+14:        $try\_time\leftarrow try\_time+1$
+15:     end while
+[…]
+17:  end if
 Figure 5: Example illustration of (1) clear match, (2) clear non-match, and (3) unclear cases.
 Preliminary Screener
-The evaluation process begins with a preliminary screener that categorizes query-code pairs into three distinct labels: (1) clear match, (2) clear non-match, and (3) unclear, as illustrated in Figure 3.
+The evaluation process begins with a preliminary screener that categorizes query-code pairs into three distinct labels: (1) clear match, (2) clear non-match, and (3) unclear, as illustrated in Fig.
+5.
 Guided by the prompt instructions, the model assesses whether the code fully addresses the query’s functional requirements.
@@
 The tests incorporate assert-based validation mechanisms targeting edge cases, functional requirements, and behavioral constraints specified in the query.
-Considering that a single query may have multiple valid explanations, the test program is generated based on one of the plausible interpretations of the query.
+Considering that a single query may have multiple valid explanations, we generate the test program based on one of the plausible interpretations of the query.
 This approach transforms ambiguous cases into executable specifications for objective validation.
@@
 This assessment is based on both the test program produced and the outcomes of its execution.
-This methodology shifts verification from theoretical code inspection to empirical observation, empowering the agent to establish evidence-based conclusions by systematically evaluating actual runtime behavior.
-By grounding judgments in dynamic test execution rather than relying exclusively on static structural analysis and semantic analysis.
-An example is provided in Appendix BB.
-An example of Final Arbiter.
-Benchmark Statistics
-In total, 412,080 labeled query-code pairs are retained. Table 1 shows the statistics of our CoSQA+ benchmark.
-Number of pairs Number of code Code average length Average test cases Executable rate
+This methodology shifts verification from theoretical code inspection to empirical observation, empowering the agent to establish evidence-based conclusions by systematically evaluating actual runtime behavior rather than relying exclusively on static structural analysis and semantic analysis.
+Fig.
+6 illustrates the process of handling an unclear case.
+Figure 6: An example for handling an unclear case through test program generation, test execution, and final arbitration.
+IV-D Benchmark Statistics
+In total, 412,080 labeled query-code pairs are retained. Table II shows the statistics of our CoSQA+ benchmark.
+TABLE II: Statistics of CoSQA+.
+"#" denotes the count, "Code avg.
+len." is the average number of characters per code snippet, and "Exec.
+rate" represents the percentage of test programs that executed successfully.
+# of pairs # of code Code avg. len. Avg. test cases Exec. rate
 412,080 132,952 378.052 4.31 83.67%
-Table 1: Statistics of CoSQA+.
-Cost Analysis
+IV-E Cost Analysis
 For data annotation, a total of 412,080 data samples are labeled by Deepseek-V3 at a cost of approximately \$257, resulting in an average cost of \$0.00062 per sample.
@@
 However, the annotation process benefits from full concurrency across multiple queries, enabling rapid completion when computational resources are sufficient.
-Evaluation
+In contrast, manual annotation by Python experts involves substantial time investment.
+On average, a Python expert spends 8 minutes per sample due to extensive research and analysis.
+For test-driven annotation, where experts design tailored test programs for code-query pairs, the time increases to 34 minutes per sample, with complex functions exceeding one hour.
+IV-F Informed consent
+This study involved human participants.
+Informed consent was obtained from all participants prior to their involvement in the study.
+Participants were informed about the purpose of the research, the procedures involved, the potential risks and benefits, and their right to withdraw from the study at any time without penalty.
+All data collected were anonymized to ensure participant confidentiality.
+V Evaluation
 In this section, we conduct experiments with corresponding analysis to answer the following research questions (RQs):
@@
 RQ3: How effective are the existing approaches in multi-choice code search on our benchmark?
-RQ1: Quality Comparison between CoSQA+ and CoSQA
+RQ4: How does each agent contribute to the performance of test-driven agents?
+V-A RQ1: Quality Comparison between CoSQA+ and CoSQA
 We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA.
 Metrics
-Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG  (Wang et al., 2013) and MRR  (Feng et al., 2020).
+Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG [14] and MRR [12].
 NDCG measures the quality of ranking by considering both the relevance and position of retrieved results.
@@
 In the first experiment, we randomly select 1000 query-code pairs labeled "1" from CoSQA+ and CoSQA, ensuring the queries are identical.
-For CoSQA+, a query may have multiple correct codes, so we randomly select one that differs from the corresponding query’s code in CoSQA.
+For CoSQA+, a query may have multiple correct codes, from which we randomly select one that differs from the corresponding query’s code in CoSQA.
 These queries are then presented to experienced programmer volunteers, who are asked to choose the code that satisfies more demand of the query.
@@
 The CoSQA+ dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division.
+TABLE III: Performance of code search models fine-tuned with different data. The test set is CSN Python.
 Model Metric Fine-tune Dataset
@@
 embedding NDCG@10 0.806 0.943 0.951
-Table 2: Performance of code search models fine-tuned with different data. The test set is CSN Python.
 Results and Analysis
-After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 73.8% of the better codes are from CoSQA+, while 26.2% are from CoSQA.
+After the volunteers compare the two datasets, we come to the result that out of the 2000 selected queries, 60.3% of the better codes are from CoSQA+, while 39.7% are from CoSQA.
 This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes.
-As shown in Table 2, the three models fine-tuned with CoSQA+ outperforms those fine-tuned with CoSQA, which indicates better training performance of CoSQA+.
+As shown in Table III, the three models fine-tuned with CoSQA+ outperform those fine-tuned with CoSQA, which indicates better training performance of CoSQA+.
 Notably, the improvement in NDCG@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
 RQ1 Summary: CoSQA+ has better quality than CoSQA. Fine-tuning CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ yields better performance than with CoSQA, especially for multiple relevant codes.
-RQ2: Performance of Agent-based Annotator
+V-B RQ2: Performance of Agent-based Annotator
 We design this experiment to assess whether our agent-based annotation approach can achieve comparable or superior quality to traditional human annotation and single LLM annotation.
 Experimental Setup
-To establish a reliable ground truth, we first have three Python experts design test programs for 300 randomly selected query-code pairs and annotate them based on test execution results.
+To establish a reliable ground truth, we first have three Python experts design test programs for 400 randomly selected query-code pairs and annotate them based on test execution results.
 The final ground truth labels are determined by majority voting among these three experts.
 We then evaluate three different annotation approaches against this ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet).
-For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators.
+For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators [61].
 Results and Analysis
+TABLE IV: Accuracy comparison of different annotation methods against ground truth.
 Method Accuracy
-Agent-based Annotator 0.964
-Human Wo. Test Program 0.890
-Deepseek-V3 0.820
-O1-mini 0.816
-Claude-3.5-Sonnet 0.830
-Table 3: Accuracy comparison of different annotation methods against ground truth.
-As shown in Table 3, our agent-based annotator achieves 96.4% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (82.0%), O1-mini (81.6%) and Claude-3.5-Sonnet (83.0%).
+Agent-based Annotator 0.920
+Human Wo. Test Program 0.880
+Deepseek-V3 0.865
+O1-mini 0.833
+Claude-3.5-Sonnet 0.845
+As shown in Table IV, our agent-based annotator achieves 92.0% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (84.0%), O1-mini (82.0%) and Claude-3.5-Sonnet (84.0%).
 This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification.
@@
 The superior performance of our agent-based annotators can be attributed to two key factors: 1) The test program provides concrete evidence of whether the code satisfies the query requirements, reducing ambiguity in the annotation process; 2) The systematic verification through test execution helps eliminate human bias and inconsistency that often occur in manual annotation.
-In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.85, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
-RQ2 Summary: Our agent-based annotator achieves 96.4% accuracy, outperforming both traditional human annotation and single LLM approaches.
+In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.818, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
+RQ2 Summary: Our agent-based annotator achieves 92.0% accuracy, outperforming both traditional human annotation and single LLM approaches.
 The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
-RQ3: Existing Methods Performance on CoSQA+
+V-C RQ3: Existing Methods Performance on CoSQA+
 This experiment aims to evaluate the effectiveness of different models in accurately retrieving code snippets that match natural language queries on the CoSQA+ benchmark.
@@
 Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly.
+To further comprehensively evaluate the performance of the models, we also introduce Mean Average Precision (MAP) and Recall as additional metrics.
+MAP provides a measure of the overall ranking quality by averaging the precision scores across all relevant results, which is especially useful for scenarios where the order of multiple correct matches matters.
+Meanwhile, Recall evaluates the model’s ability to retrieve all relevant code snippets from the dataset, ensuring that no important results are missed.
+Together, these metrics offer a more robust assessment of the model’s effectiveness in semantic code search tasks.
 The tested methods include both traditional information retrieval techniques and advanced deep learning-based models.
-Specifically, we compare a traditional Bag of Words (BoW) approach (Zhang et al., 2010) with state-of-the-art semantic code search models, namely: CodeBERT (Feng et al., 2020), UniXcoder (Guo et al., 2022), and CodeT5+ embedding (Wang et al., 2023).
-In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3  (Sturua et al., 2024), All-MiniLM-L12-v2 (Reimers and Gurevych, 2021), All-mpnet-base-v2  (Song et al., 2021), and Multilingual-e5-large  (mul, 2023).
+Specifically, we compare a traditional Bag of Words (BoW) approach [27] with state-of-the-art semantic code search models, namely: CodeBERT [12], UniXcoder [31], and CodeT5+ embedding [34].
+In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3 [35], All-MiniLM-L12-v2 [40], All-mpnet-base-v2 [41], and Multilingual-e5-large [38].
 Results and Analysis
-Method NDCG@10 MRR
-BoW 0.046 0.034
-CodeBERT 0.232 0.175
-UniXcoder 0.187 0.140
-CodeT5+ embedding 0.179 0.136
-Jina-embedding-v3 0.225 0.171
-All-MiniLM-L12-v2 0.214 0.163
-All-mpnet-base-v2 0.227 0.172
-Multilingual-e5-large 0.218 0.166
-Table 4: Performance of Various Methods on CoSQA+ Using NDCG@10 and MRR Metrics.
-While traditional information retrieval methods, such as BoW, are computationally efficient, Table 4 demonstrates that they struggle to capture the semantic subtleties needed to accurately map natural language queries to relevant code snippets, particularly when multiple valid solutions exist.
-Among the deep learning-based models, CodeBERT achieves the highest performance with an NDCG@10 of 0.232 and an MRR of 0.175.
-This strong performance can be attributed to its encoder-only architecture and extensive pre-training on large-scale paired natural and programming language corpora, especially Python, which enhances its code understanding capabilities.
+TABLE V: Performance of Various Methods on CoSQA+ Using NDCG@10, MRR, MAP and Recall Metrics.
+Method NDCG@10 MRR MAP Recall
+BoW 0.037 0.027 0.027 0.016
+CodeBERT 0.232 0.175 0.164 0.149
+UniXcoder 0.187 0.140 0.132 0.110
+CodeT5+ embedding 0.179 0.136 0.127 0.101
+Jina-embedding-v3 0.225 0.171 0.160 0.145
+All-MiniLM-L12-v2 0.214 0.163 0.153 0.135
+All-mpnet-base-v2 0.227 0.172 0.162 0.145
+Multilingual-e5-large 0.218 0.166 0.156 0.136
+While traditional information retrieval methods, such as BoW, are computationally efficient, Table V demonstrates that they struggle to capture the semantic subtleties needed to accurately map natural language queries to relevant code snippets, particularly when multiple valid solutions exist.
+Among the deep learning-based models, CodeBERT achieves the highest performance with NDCG@10 of 0.232, MRR of 0.175, MAP of 0.164 and Recall of 0.149.
 Other embedding-based models also show substantial improvements over traditional methods, highlighting the benefits of modern deep learning architectures for semantic code search.
+This strong performance can be attributed to their encoder-only architecture and extensive pre-training on large-scale paired natural and programming language corpora, especially Python, which enhances their code understanding capabilities.
+The embedding models originate from different time periods and vary in parameter sizes.
+Surprisingly, CodeBERT outperforms the latest embedding models, which may be attributed to the fact that CodeBERT tested here is fine-tuned specifically for code search tasks, whereas the latest embedding models are designed for general-purpose tasks.
+Some smaller-parameter embedding models demonstrate impressive performance.
+For instance, all-mpnet-base-v2, with only 109M parameters, and all-MiniLM-L12-v2, with a mere 33M parameters, still exhibit remarkable performance, surpassing many models with larger parameter sizes.
+This represents a significant advancement in embedding models in recent years, providing a solid foundation for the practical implementation of semantic search based on large models.
+From the evaluation results, the performance differences among embedding models seem not significant.
+However, due to the richness of CoSQA+, which effectively eliminates random factors, even subtle differences in metrics can reflect the overall capability gap between models.
+Moreover, embedding models generally perform poorly on our benchmark.
+From the perspective of model design, embedding models are mainly used for retrieval tasks.
+To improve inference speed, their weight parameters are usually small, limiting their ability to deeply understand query sentences and code.
+The CoSQA+ is challenging in two main aspects: First, the queries originate from CoSQA, which are real user queries filtered from the Bing search engine.
+These queries are often expressed casually and may contain spelling or grammar errors.
+Second, the codebase used for retrieval is not only large in scale but also contains many confusing code snippets.
 RQ3 Summary: The evaluation on CoSQA+ clearly indicates that deep learning-based models significantly outperform traditional retrieval methods for semantic code search.
 Notably, CodeBERT leads the performance ranking by achieving the highest scores in both NDCG@10 and MRR, while the alternative embedding-based models also offer competitive results.
-Conclusion
-In this paper, we present CoSQA+, the first large-scale multi-choice code search dataset with 412,080 pairs.
-Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks.
-Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA.
-The test-driven annotation approach achieves a high accuracy of 96.4%, outperforming traditional human annotation and single LLM methods.
-Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks.
-We believe our agents-annotated CoSQA+ benchmark will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task
-Ethical consideration
+V-D RQ4: Contributions of Different Components in Test-Driven Agents
+This experiment is to assess the contribution of each component in our test-driven agent framework.
+Experimental setup
+The final test-driven agents comprise the following components: Preliminary Screener (PS), Test Program Generator (TPG), Test Executor (TE), Bug Fixer (BF), and Final Arbiter (FA).
+We divided the agent into different configurations by combining these components in various ways.
+For each configuration, we compared the predicted labels with the ground truth labels (obtained from human evaluation based on the test programs) to calculate the accuracy.
+Additionally, we compare the Simulated Executor with both the Test Executor and the Bug Fixer to evaluate whether the large language model can accurately simulate execution without actual interpretation.
+Unlike the Test Executor, which runs the test program using a Python interpreter, the Simulated Executor leverages the LLM’s reasoning capabilities to emulate execution and directly predict the outcome.
+Result and Analysis
+Table VI shows the accuracy for different configurations.
+Configuration Accuracy
+Single LLM 0.865
+PS + FA 0.840
+PS + TPG + FA 0.833
+PS + TPG + TE + FA 0.835
+PS + TPG + Simulated Executor + FA 0.840
+PS + TPG + TE + BF + FA 0.920
+TABLE VI: Accuracy of Various Test-Driven Agent Configurations (Abbreviated)
+The results indicate that the baseline Single LLM achieves an accuracy of 0.865.
+When the system uses only the Preliminary Screener in combination with the Final Arbiter, the accuracy slightly drops to 0.840.
+Incorporating the Test Program Generator alongside these yields an accuracy of 0.833, suggesting that generating test programs without execution may introduce noise or irrelevant information, which can harm performance [62].
+Adding the Test Executor to the configuration improves the performance marginally (0.835), and using a Simulated Test Executor does not provide a significant advantage (0.840).
+This aligns with the observation that simulated execution cannot fully capture the complexity of the real compile-and-run process, potentially introducing inaccuracies or misleading feedback.
+Notably, the inclusion of the Bug Fixer increases the accuracy substantially to 0.920.
+This enhancement demonstrates that the Bug Fixer is crucial for processing the test programs effectively, as it enables the system to generate real execution results that are vital for improving prediction accuracy.
+RQ4 Summary: The Bug Fixer is the most important component in the test-driven agent framework as it makes the test program executable and yields effective test results.
+However, it relies on the preceding steps, such as generating the test program.
+Simulated execution results cannot replace the benefits of real execution feedback.
+VI Discussion
+VI-A Test Misjudgment
+Previous studies show that as the size of the codebase increases, the challenges faced by embedding models also increase [63, 64].
+This conclusion aligns with intuition, but with the analysis of other datasets and the construction of CoSQA+, we find that the decline in performance is not only due to the need to exclude more confusing codes but also because of the increased likelihood of test misjudgments.
+For example, a LLM may select a piece of code that meets the query requirements, but this query-code pair is not included in the candidate code set.
+As a result, this correct selection is misjudged as incorrect.
+When constructing code search datasets, query sets and code sets are typically extracted from original query-code pairs for test.
+However, the number of all possible query-code pairs (similar to a Cartesian product) formed between the query set and the code set far exceeds the number of original pairs.
+Existing methods assume that the code of a pair outside the original pairs cannot correctly match the query, but this assumption is not entirely valid.
+As shown in Fig.
+7, in CoSQA+, excluding codes without satisfied query, 36.7% of codes can satisfy two or more queries.
+Constructing a dataset completely free of misjudgments is almost impossible, as it would require checking all possible pairs.
+As can be seen by combining Fig 3 and Fig.
+7, the approach adopted in CoSQA+, where 20 candidate codes are annotated for each query, helps mitigate these testing misjudgments.
+Another possible measure is to invoke test-driven annotation agents when a code search method selects a previously unseen query-code pair to determine whether the selected code matches the query.
+However, this approach may introduce significant computational overhead, especially when the candidate code repository is very large.
+Figure 7: Distribution of codes by the number of queries they satisfy
+VI-B Inherent Ambiguity in Query and Code Labels
+A major challenge in code search is the inherent ambiguity in queries and code.
+Queries are often vague or poorly phrased, leading to multiple interpretations, while code snippets may be incomplete or abstracted, making correctness subjective.
+For instance, a query like "how to check if a folder exists and create it if not" can be interpreted differently based on assumptions about paths or edge cases.
+Similarly, code may solve the problem partially or in a non-standard way, complicating strict labeling ("exact match").
+Test-driven annotation helps reduce ambiguity by generating executable test programs aligned with the query’s intent.
+However, variability in query interpretation and incomplete code still introduce labeling challenges.
+Future improvements will require clearer queries and more complete code snippets to minimize ambiguity.
+VI-C Ethical Consideration
 In this work, we prioritize ethical considerations by the responsible use of data and methodologies.
@@
 We encourage the community to adopt ethical guidelines when using our benchmark and tools, ensuring that their application aligns with positive social impact.
-Limitations
-We are committed to using LLM-based test-driven annotation to construct high-quality function-level multi-choice code search datasets, but there are still several limitations.
-Firstly, due to resource constraints, we only apply the construction method to Python code.
-The method is generalizable, and we plan to extend it to other programming languages.
-Secondly, we conduct extensive experiments, research, and manual validation to demonstrate the effectiveness of the construction method and the quality of the dataset, but due to limitations in time and resources, CoSQA+ may contain some biases or ignore some scenarios.
-Finally, although various mechanisms are designed to improve the success rate and accuracy of automated annotation and achieve promising results, factors such as ambiguity in query statements, complexity of code functionality, and uncertainty in program execution environments prevent ideal outcomes for few query-code pair annotations.
-We plan to continue exploring and improving the framework and methods for automated annotation in the future.
-Appendix A A. Prompts for Test-Driven Agents
-Preliminary Screener
-1def get_preliminary_screening_prompt( query, function_code):
-2 prompt = f"""
-3 Instructions for Evaluation:
-4 You are provided with a query that specifies requirements for a function and a block of function code.
-5 Your goal is to determine if the code fully matches the query’s requirements.
-6 - If the code clearly does not satisfy the query, output preliminary_screening: 0 and provide a corresponding explanation.
-7 - If the code clearly satisfies the query, output preliminary_screening: 1 and provide a corresponding explanation.
-8 - If the determination is ambiguous or if further testing using a test_program_generation agent is needed, output preliminary_screening: 0.5 and provide a corresponding explanation.
-9
-10 Ensure your answer is formatted exactly as follows:
-11 preliminary_screening: [0/0.5/1],
-12 explanation: <Your explanation in English>
-13
-14
-[…]
-[…]
-17 """
-18 return prompt
-Test Program Generator
-1def get_test_program_prompt(query, function_code):
-2 prompt = f"""
-3 Task Description:
-4 Please generate a test program based on the provided query and function code. The test program should:
-5 1. Contain assert statements in the main function to verify that the function code correctly implements the behavior described in the query.
-6 2. **Do not** test other classes or functions generated, only focus on testing the behavior specified in the query.
-7 3. The generated test program should be complete, executable, and must test the functionality described by the query using assert statements.
-8
-9 For example, given the following query and function code:
-10
-11 query: "python remove all empty items in list"
-12 code:
-13 def remove_from_list(list, remove_items):
-14 list = return_list(list)
-15 remove_items = return_list(remove_items)
-16 for remove in remove_items:
-17 if remove in list:
-18 list.remove(remove)
-19 return list
-20
-21 The generated program should:
-22 Test the behavior described in the query (removing empty items from a list).
-23 Include only the assert statements in the main function.
-24 Be a fully executable test program that can verify the behavior of function code.
-25
-26 Example Test Program Output:
-27
-28 # Additional function definition
-29 def return_list(input_list):
-30 return [item for item in input_list if item]
-31
-32 # Original function code
-33 def remove_from_list(list, remove_items):
-34 list = return_list(list)
-35 remove_items = return_list(remove_items)
-36 for remove in remove_items:
-37 if remove in list:
-38 list.remove(remove)
-39 return list
-40
-41 # Test code
-42 def main():
-43 # Test case 1: Remove empty items from the list
-44 list1 = [1, 2, "", 4, None, "test", ""]
-45 remove_items1 = [""]
-46 result1 = remove_from_list(list1, remove_items1)
-47 # Assert that empty strings have been removed from the list
-48 assert result1 == [1, 2, 4, None, "test"]
-49
-50 # Test case 2: Remove None from the list
-51 list2 = [1, 2, "", 4, None, "test", ""]
-52 remove_items2 = [None]
-53 result2 = remove_from_list(list2, remove_items2)
-54 # Assert that None has been removed from the list
-55 assert result2 == [1, 2, "", 4, "test", ""]
-56
-57 # Test case 3: Ensure no changes when removing an item not in the list
-58 list3 = [1, 2, 3]
-59 remove_items3 = [4]
-60 result3 = remove_from_list(list3, remove_items3)
-61 # Assert that the list remains unchanged
-62 assert result3 == [1, 2, 3]
-63
-64 # Test case 4: Empty list case
-65 list4 = []
-66 remove_items4 = [""]
-67 result4 = remove_from_list(list4, remove_items4)
-68 # Assert that the result is still an empty list
-69 assert result4 == []
-70
-71 print("All tests passed.")
-72
-73 if __name__ == "__main__":
-74 main()
-75
-76 Now please directly and only generate the complete test program based on the instructions above. Do NOT add any explanation.
-77 query: {query}
-78 code: {function_code}
-79 """
-80
-81 return prompt
-Bug Fixer
-1def get_pip_install_command_prompt( test_output: str, test_program: str) -> str:
-2 prompt=f"""You are a Python development assistant. The user encountered a ModuleNotFoundError when running their test code. Here are the details:
-3 Test Output:
-4 {test_output}
-5
-6 Test Program:
-7 {test_program}
-8
-9 Please determine the correct pip install command(s) required to install the missing module(s).
-Consider that the PyPI package name might differ from the module name.
-Provide only the necessary pip install command(s), each on a separate line.
-Do not include any other text or explanation.
-10
-11 Examples of valid responses:
-12 1. If missing ’cv2’:
-13 pip install opencv-python
-14
-15 2. If missing ’yaml’:
-16 pip install pyyaml
-17
-18 3. If multiple packages needed:
-19 pip install numpy
-20 pip install pandas
-21
-22 Now, provide the correct pip install command(s) for the current error:
-23 """
-24 return prompt
-Final Arbiter
-1 def get_final_judge_prompt( function_code: str, query: str, test_program: str, test_result: str,
-2 test_output: str) -> str:
-3 prompt = f"""
-4 You are an code expert in judging whether a function code can fully satisfy the requirement of the query. You can refer to a generated test program and its test result and test output.
-5
-6 You are provided with the following inputs:
-7
-8 1. Query:
-9 {query}
-10
-11 2. Function Code:
-12 {function_code}
-13
-14 3. Test Program:
-15 {test_program}
-16
-17 4. Test Result:
-18 {test_result}
-19
-20 5. Test Output:
-21 {test_output}
-22
-23 Your Task:
-24 Evaluate whether the provided function code completely satisfies the query’s requirements based on a detailed analysis of the function code and the results and the output from running the test program.
-25
-26 Judgment Guidelines:
-27 - If the function code clearly fails to meet the query requirements (supported by incorrect logic or failing test results), then output:
-28 final_verdict: 0,
-29 final_explanation: "A clear explanation describing why the function code does not meet the query requirements, referencing any issues in the code and test failures."
-30
-31 - If the function code fully meets all the query requirements (supported by correct logic and successful test results), then output:
-32 final_verdict: 1,
-33 final_explanation: "A clear explanation describing how the function code meets all the query requirements, referencing the successful test outcomes and correct implementation."
-34
-35 Important:
-36 - Your final output must strictly adhere to the following format:
-37 final_verdict: [0 or 1],
-38 final_explanation: <Your explanation in English>
-39 - Only output 0 or 1 as the final verdict.
-40
-41 Please provide your final judgment as instructed.
-42 """
-43 return prompt
-B. An example of Final Arbiter
-This figure illustrates the process of handling an unclear case.
-The test program generator generates a program, which is then executed by the test executor.
-The test results are returned to the final arbiter, which assigns a label to the case.
-Figure 4: Process of Handling an Unclear Case through Test Program Generation, Execution, and Final Arbitration
-C. Code Search Behavior Survey
-We conduct this survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.
-Survey Demographics and Methodology
-A total of 102 participants complete the survey, predominantly undergraduate students (89.22%) with some master’s students (8.82%).
-The majority of participants (92.16%) have formal computer science education, and most have 1-5 years of programming experience (66.67%), while 30.39% have less than one year of experience.
-All participants are Chinese university students who use Python for development, with 52.94% considering Python as their primary programming language.
-The survey is conducted anonymously through an online platform, with all questions presented in Chinese.
-No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
-Survey Results and Analysis
-Our survey reveal several significant patterns in code search behavior.
-Most notably, when 66.67% typically needing 2-3 code examples to understand and implement a solution.
-Participants evaluate search results based on several key factors, primarily focusing on implementation relevance (65.69%) and solution differences (56.86%).
-When presented with multiple code examples, they particularly value the clarity of implementation differences and the completeness of edge case handling.
-Survey Questions
-Table 5 shows the key questions from our survey (translated from Chinese).
-Questions Response Rate (%)
-A B C D
-How often do you perform code searches during programming-related tasks per day?
-(A: >10 times, B: 6-10, C: 3-5, D: 1-2) 50.00 19.61 14.71 15.68
-In your daily development, how many code examples do you typically need to reference for a query (such as API usage, algorithm implementation)
-(A: Only 1, B: 2-3, C: 4-5, D: >5) 16.67 66.67 12.75 3.92
-In which scenarios do you most frequently need to reference multiple code examples when performing code searches? [Multiple Choice]
-(A: API usage, B: Algorithm implementation, C: Performance optimization solutions, D: Other) 68.63 60.78 51.96 47.06
-When performing code searches, which better matches your habits?
-(A: Only need to find one code snippet that meets the requirements, B: Need to find more than one snippet that meets the requirements) 49.02 50.98 - -
-When search returns multiple code examples, which aspects do you focus on most? [Multiple Choice]
-(A: Differences between code examples, e.g.
-comparison between different implementation methods, B: Relevance of each example, e.g.
-degree of matching with the query, C: Whether the display order of examples meets expectations, e.g.
-most relevant first, D: Whether examples cover common use cases or edge conditions) 56.86 65.69 49.02 38.24
-When evaluating code search results, which do you prioritize more?
-(A: Position of the first result that meets requirements, B: Average position of the first few results that meet requirements, C: Diversity of results, D: Other) 44.12 37.25 18.63 0.00
-When searching for code examples, which type of function structure do you prefer?
-(A: Functions with both input and output, B: Functions with input but no output, C: Functions with output but no input , D: Functions with neither input nor output) 71.57 16.67 8.82 2.94
-When searching for code examples, which type of function structure do you end up using?
-(A: Functions with both input and output, B: Functions with input but no output, C: Functions with output but no input , D: Functions with neither input nor output) 62.75 22.55 11.76 2.94
-Table 5: Distribution of survey responses (n=102). For each question, the options are listed in parentheses.
-D. Distribution of Queries Based on Exactly Matched Codes
-The table 6 illustrates the distribution of queries according to the number of exactly matched codes, providing insight into the frequency of queries with varying code match counts.
-Table 6: Distribution of Queries Based on Exactly Matched Codes
-Number of Exactly Matched Codes (n) Query Count
-1 2907
-2 2445
-3 2121
-4 1748
-5 1429
-6 1266
-7 1108
-8 873
-9 724
-10 628
-11 481
-12 398
-13 345
-14 220
-15 155
-16 126
-17 78
-18 44
-19 19
-20 3
+VII Conclusion
+In this paper, we present CoSQA+, the first large-scale multi-choice code search dataset with 412,080 pairs.
+Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks.
+Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA.
+The test-driven annotation approach achieves a high accuracy of 92.0%, outperforming traditional human annotation and single LLM methods.
+Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks.
+We believe our agents-annotated CoSQA+ benchmark will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task.

### v6

@@
 We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system.
-Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 92.0%.
+Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 93.9%.
 Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA.
 Models trained on CoSQA+ exhibit improved performance.
-We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
-Index Terms:
-software engineering, information search and retrieval, human-computer interaction.
+We publicly release both CoSQA+_all, which contains 412,080 agent-annotated pairs, and CoSQA+_verified, which contains 1,000 human-verified pairs, at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
 11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
@@
 Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems [6].
-Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories.
-Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR) [7, 8, 9, 10, 11, 12].
-According to our survey of 102 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 8.31 times per programming day on average.
-For each query, they typically reference 2.64 code examples to find the necessary information.
-These findings highlight a key need in creating a multi-choice code search benchmark: effectively pairing real-world queries with multiple correct and relevant code snippets.
-However, existing approaches face significant challenges in scalability, annotation quality, and evaluation metrics:
+For the given query, while existing semantic code search studies focus on retrieving relevant codes, we define a different task: to retrieve functionally matching codes that satisfy the query’s requirements when executed in a complete context.
+Although one query can be matched by multiple codes, existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR) [7, 8, 9, 10, 11, 12].
+This design leads to a critical mismatch with real-world practices.
+According to our survey of 200 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 7.79 times per programming day on average.
+For each query, they typically reference 2.83 code examples to find the necessary information.
+Furthermore, participants report that 63.2% of their queries commonly result in multiple valid code snippets.
+This reveals that developers operate in a multi-choice context, yet existing benchmarks adhere to a single-choice evaluation paradigm with only one "correct" code snippet per query.
+The construction of benchmarks that reflect this multi-choice reality faces significant challenges, as methodologies designed for the prevalent one-to-one task (e.g., CoSQA [8], CodeSearchNet [7]) are hampered by issues of scalability, accuracy, and evaluation:
   1. C1
@@
 Both human and LLM-based annotation approaches suffer from accuracy issues.
-Human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations.
+Human annotators typically assess code matching based on surface-level comprehension rather than verifying functional correctness [13] through test execution, which may lead to inaccurate annotations due to a lack of full understanding of syntax, semantics, and behavioral correctness.
 While Li et al.
@@
   3. C3
-    Inadequate Evaluation Metrics for Code Search.
-Most code search benchmarks employ metrics like MRR or Pass@k [10] which primarily evaluate single-result retrieval performance.
-This overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations.
-While alternative metrics like MAP and Recall@k [7] exist for evaluating multiple code matches, they underemphasize the importance of top-ranked results.
-This is particularly problematic as our survey reveals users heavily rely on first-page search results, highlighting the need for better evaluation metrics.
-To address these limitations, we present a new benchmark CoSQA+, with 412,080 pairs of web queries from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search.
+    Lack of Evaluation in Multi-Choice Scenarios.
+One-to-one Conventional one-to-one matching benchmarks typically rely on MRR, which measures the rank position of the first relevant code snippet.
+However, in practical scenarios, developers often seek multiple code examples for reference, where the key concern is how many relevant snippets appear within the top-10 results or on the first page.
+MRR and other single-choice benchmarking approaches fail to capture this dimension of performance.
+To address these limitations, we present a new benchmark CoSQA+ , comprising two releases: CoSQA+_all with 412,080 pairs of web queries from CoSQA and repository code curated from CodeSearchNet, and CoSQA+_verified with 1,000 human-verified pairs that serve as a gold-standard evaluation subset.
+The benchmark is specifically designed for multi-choice code search and adopts MAP as the primary evaluation metric.
 To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a preliminary screener, a test program generator, a test executor, a bug fixer, and a final arbiter.
 The label indicates whether each code snippet exactly matches its corresponding query requirements.
-Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks [14].
-We use it to measure 8 code search methods, 7 of which are embedding-based.
-To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 400 randomly selected query-code pairs from CoSQA+.
+Specifically, exactly matched codes refer to codes whose functionality fully satisfies or exceeds the requirements specified in the query.
+Furthermore, we evaluate 8 code search methods—7 of which are embedding-based.
+To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 1000 randomly selected query-code pairs from CoSQA+.
 Three Python experts first design test programs for these pairs to establish ground truth annotations.
 We then compare the accuracies among test-driven agents, python expert annotators without tests, and a single LLM.
-Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 92.0%.
+Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 93.9%.
 Notably, 83.67% of the test programs automatically generated by test-driven agents are executable.
 Furthermore, using identical data sources (queries from CoSQA and code from CodeSearchNet), we evaluate CoSQA+ against CoSQA as fine-tuning corpora for three models: CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings.
-Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both NDCG@10 and MRR metrics.
+Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both MAP(Mean Average Precision)@10 and MRR metrics.
+Furthermore, we demonstrate that the test-driven agent generalizes beyond Python, achieving consistent accuracy gains of 10.33% on average across PHP, Java, and Go (RQ5).
 Our contributions can be summarized as follows:
-  • To the best of our knowledge, we develop the first Multiagent annotator with 92.0% accuracy for code search dataset.
+  • To the best of our knowledge, we develop the first Multiagent annotator with 93.9% accuracy for code search dataset.
   • To the best of our knowledge, we construct CoSQA+ which is the first benchmark designed for evaluating multi-choice code search.
   • We enhance the fully automated construction process through candidate pair construction by multiple models and annotation by agents.
-  • We publicly release CoSQA+ to facilitate further research in semantic code search.
+  • We publicly release both CoSQA+_all and CoSQA+_verified to facilitate further research in semantic code search.
 II Motivation Survey
 Do developers really need multi-choice code search?
-To answer this question, we conduct a survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.Fig.
-1 provides a multi-choice code search example
+To answer this question, we conduct a survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.
+Fig.
+1 provides a multi-choice code search example.
+This example illustrates a common real-world scenario where query ambiguity leads to multiple legitimate interpretations.
+The term "illegal character" can reasonably refer to filesystem-invalid characters, non-ASCII characters, escape sequences, or special symbols, each representing valid use cases in different programming contexts.
 Figure 1: An example of multi-choice code search. The query is vague in its definition of "illegal character," allowing eight candidate code snippets to match under different interpretations.
 Methodology
-The survey targets 102 participants, predominantly undergraduate students (89.22%) with a smaller proportion of master’s students (8.82%).
-Most of the participants (92.16%) have formal computer science education, and their experience in Python programming varies.
-The majority (66.67%) have between 1 and 5 years of programming experience, while 30.39% have less than one year of experience.
-The participants are selected from programmers using Python as their primary programming language (52.94%).
-The survey is conducted anonymously through an online platform.
+We recruited participants through online anonymous questionnaire platforms, targeting two main groups: current university students and working technical personnel in computer science or related fields.
+The survey collected responses from 200 participants, including 148 current students and 52 working professionals.
+In terms of educational background, the majority of participants hold bachelor’s degrees (87.5%) with a smaller proportion holding master’s or doctoral degrees (11.5%).
+Their Python programming experience varies, with the majority (59.0%) having between 1 and 5 years of programming experience, while 16.0% have between 5 and 10 years of experience.
+Among all participants, 64.0% use Python as their primary programming language.
 No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
 Results and Analysis
-Our survey reveals several significant patterns in code search behavior.
-Most notably, when 66.67% typically needing 2-3 code examples to understand and implement a solution.
-This suggests that the current code search task, which typically prioritizes a single result, may need to evolve to provide multiple relevant code snippets.
-Participants evaluate search results based on several key factors, primarily focusing on implementation relevance (65.69%) and solution differences (56.86%).
+Our survey reveals several significant patterns in code search behavior across different user groups.
+Importantly, the need for multiple code examples is not limited to undergraduate students but extends to experienced developers as well.
+We find that 75.0% of working technical professionals typically require 2-3 code examples to understand and implement a solution, compared to 63.5% of current students.
+This finding challenges the assumption that multi-choice code search primarily serves educational purposes for beginners.
+In fact, experienced developers demonstrate an even higher tendency to reference multiple implementations, particularly when dealing with complex or unfamiliar tasks.
+Additionally, the participants reported that 63.2% of their queries commonly result in multiple code snippets that meet their requirements, indicating that query ambiguity leading to multiple valid solutions is a prevalent phenomenon in real-world development scenarios.
+Code searchers prefer function codes with clear input and output (85.5%), indicating a strong preference for code examples that explicitly define parameters and expected results regardless of their experience level.
+Participants evaluate search results based on several key factors, primarily focusing on solution differences (66.5%) and implementation relevance (66.5%).
 These preferences emphasize that developers do not just want the closest match to their query, but also value diverse approaches that help them understand the problem and solution better.
@@
 For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets [57].
-(2) Through rigorous filtering and validation processes, the queries demonstrate strong consistency in format and clear programming intent.
+(2) Through heuristic filtering and manual annotation, the queries maintain authentic code search intent while reducing non-code-search-related noise.
 This careful curation helps maintain high data quality and reduces noise in the training process [58].
@@
 Code Source
-To the best of our knowledge, CodeSearchNet is the only origin dataset specifically designed for Python functions in the domain of code search.
-While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet.
+To the best of our knowledge, CodeSearchNet is the only large-scale, real-world Python function dataset explicitly designed for code search.
+While other benchmarks (e.g., CoSQA [8], FunSearch [59], AdvTest [10]) include Python code, their codebases are derived from CSN — either by reusing its functions directly or by further processing its splits.
+This makes CSN the de facto origin for Python code-search evaluation.
 The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages [7].
-According to our survey, code searchers prefer function codes with clear input and output.
-These codes are high quality with higher readability, testability and maintainability [59].
-To extract these high-quality codes from CodeSearchNet’s 1,156,085 Python functions, we filter out 653,994 functions using a rule-based approach.
-This involves parsing the code into an Abstract Syntax Tree (AST), identifying function nodes with at least one input parameter and a valid return statement.
-Such testable functions lay the foundation for test-driven agents to design test program and improve dataset quality.
+According to our survey, 85.5% of developers report that functions with explicitly defined inputs and outputs best meet their needs, reflecting real-world developer preferences.
+Such functions are consistently rated higher in readability, testability, and maintainability [60].
+While functions with outputs but no inputs may still be relevant and testable in some contexts, they are often considered poorly designed.
+These functions typically exhibit unclear logic, rely on implicit global states or hard-coded constants, and are difficult to reuse across different scenarios [60, 61].
+To assess data quality and relevance, we carefully reviewed 1,995 functions and found that only 16.25% of functions without input or output parameters are relevant to typical code search tasks, compared to 80.97% of functions with both input and output parameters.
+Including parameter-less functions would therefore introduce noise and degrade overall dataset quality.
+Therefore, to construct a high-quality Python code corpus from CodeSearchNet’s 1,156,085 functions, we employ a rule-based AST filter to retain only those functions having at least one input parameter and a valid return statement, yielding 653,994 candidates.
+By concentrating on these clearly specified and testable functions, our dataset preserves sufficient scale, higher quality, and provides a solid foundation for test-driven agents to design and execute meaningful test programs, ultimately improving downstream model quality and supporting accurate code search evaluation.
 IV-B Candidate Pairs Construction
@@
 Construction Steps
-The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matching codes from a large codebase.
+The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matched codes from a large codebase.
 This forms the foundation for annotating query-code pairs in the CoSQA+ dataset.
@@
 Then, these embeddings are used to calculate cosine similarity scores, averaging similarity results across the three models.
+Cosine similarity is calculated as:
+${cosine\_similarity}=\frac{\mathbf{v}_{\text{query}}\cdot\mathbf{v}_{\text{code}}}{\|\mathbf{v}_{\text{query}}\|\|\mathbf{v}_{\text{code}}\|}$
+where $\mathbf{v}_{\text{query}}$ and $\mathbf{v}_{\text{code}}$ are the embedding vectors of the query and code, respectively.
 Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query.
@@
 CodeT5+ embedding 0.088
-The relatedness between the query and code is quantified by the cosine similarity, calculated as:
-${cosine\_similarity}=\frac{\mathbf{v}_{\text{query}}\cdot\mathbf{v}_{\text{code}}}{\|\mathbf{v}_{\text{query}}\|\|\mathbf{v}_{\text{code}}\|}$
-where $\mathbf{v}_{\text{query}}$ and $\mathbf{v}_{\text{code}}$ are the embedding vectors of the query and code, respectively.
+The relatedness between the query and code is quantified by the cosine similarity.
 By taking the average similarity score from these three models, we achieve a more balanced and fair assessment of query-code pair quality.
@@
 To efficiently and accurately annotate the large number of query-code pairs, we design the test-driven annotation agents based on DeepSeek-V3 [62], using the algorithm shown in Algorithm 1.
-Initially, A preliminary screener determines whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment.
+Initially, a preliminary screener determines whether the query clearly matches or does not match the code, or if it is unclear.
+For clear cases, we directly assign labels based on the initial assessment.
 For uncertain cases, we determine the labels based on the test execution results and final judgment.
@@
 Output: label $l\in\{0,1\}$
-1:  $assessment\leftarrow\text{preliminary\_screener}(q,c)$
 […]
+2:if $assessment=\text{``clearly match''}$ then
 3:  return 1
-[…]
+4:else if $assessment=\text{``clearly non-match''}$ then
 5:  return 0
-[…]
+6:else if $assessment=\text{``unclear''}$ then
 7:  $test\_prog\leftarrow\text{generate\_test\_program}(q,c)$
@@
 9:  $try\_time\leftarrow 0$
-10:     while $exec\_res.\text{error}\ \textbf{and}{}\ try\_time\leq max\_try\_time$ do
+[…]
 11:   $terminal\_commands\leftarrow\text{fix\_test\_program}(test\_prog,exec\_res.\text{error})$
-[…]
+12:   $\text{run\_terminal}(terminal\_commands)$
 13:   $exec\_res\leftarrow\text{execute\_test}(test\_prog)$
@@
 15:  end while
-[…]
+16:  return $\text{final\_arbiter}(q,c,test\_prog,exec\_res)$
 17:end if
@@
 This approach transforms ambiguous cases into executable specifications for objective validation.
+To be clear, the mere fact that code runs does not guarantee it is relevant to the query.
+The agents, therefore, generate test programs that execute the code and compare its actual outputs or side effects with the results expected from the query’s intent.
+Execution logs and assertion outcomes are then passed to the final arbiter, which makes a holistic judgment rather than relying solely on execution success or failure.
+In this way, verification shifts from a literal understanding of the snippet to functional validation of its behaviour.
 Test Executor and Bug Fixer
@@
 RQ4: How does each agent contribute to the performance of test-driven agents?
+RQ5: Can the test-driven agent generalize across different programming languages?
 V-A RQ1: Quality Comparison between CoSQA+ and CoSQA
-We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA.
+We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA by examining how well models can learn from these datasets.
+Here, we treat CoSQA and CoSQA+ as training datasets rather than benchmark suites, focusing on their effectiveness for model training rather than their utility as evaluation benchmarks.
+Note that this dual functionality — serving as both training corpus and evaluation benchmark — follows established precedent in code search research, where datasets like CodeSearchNet [7], CoSQA [8] and StaQC [11] similarly function in both capacities.
+For this specific comparison, we focus on the training aspect to assess dataset quality and learning effectiveness.
 Metrics
-Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG [14] and MRR [12].
-NDCG measures the quality of ranking by considering both the relevance and position of retrieved results.
-For a given query $q_{i}$, we calculate NDCG@k as:
-$NDCG@k=\frac{DCG@k}{IDCG@k}$
-where DCG@k is the Discounted Cumulative Gain up to position k, and IDCG@k is the ideal DCG@k (the maximum possible DCG@k).
-The final NDCG score is averaged across all queries.
+Following previous work, we employ two metrics to evaluate the performance of code search models: MAP and MRR [12].
+MAP measures the overall ranking quality by averaging precision scores across all relevant results for each query.
+For a given query $q_{i}$ with $R_{i}$ relevant results, we calculate Average Precision (AP) as:
+$AP@k=\frac{1}{R_{i}}\sum_{j=1}^{k}P(j)\cdot rel(j)$
+where $P(j)$ is the precision at position $j$, and $rel(j)$ is an indicator function that equals 1 if the result at position $j$ is relevant.
+The final MAP score is obtained by averaging AP across all queries:
+$MAP=\frac{1}{N}\sum_{i=1}^{N}AP@k_{i}$
+where $N$ is the total number of queries. This metric is particularly suitable for code search as it properly handles cases where multiple correct results may exist for a query.
 MRR evaluates the average reciprocal of the rank at which the first relevant code appears:
@@
 Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
-For the second experiment, We fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with NDCG@10 and MRR as evaluation metrics.
+For the second experiment, We fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with MAP and MRR as evaluation metrics.
 From the CSN benchmark we use for validation, known as CSN99, we extract Python test samples and build 399 query-code pairs for 95 queries, referred to as CSN Python.
@@
 CodeBERT MRR 0.952 0.939 0.966
-NDCG@10 0.958 0.937 0.962
+MAP@10 0.935 0.900 0.938
 UniXcoder MRR 0.891 0.951 0.954
-NDCG@10 0.890 0.946 0.952
+MAP@10 0.828 0.912 0.926
 CodeT5+ MRR 0.794 0.938 0.940
-embedding NDCG@10 0.806 0.943 0.951
+embedding MAP@10 0.737 0.910 0.928
 Results and Analysis
-After the volunteers compare the two datasets, we come to the result that out of the 2000 selected queries, 60.3% of the better codes are from CoSQA+, while 39.7% are from CoSQA.
+After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 60.3% of the better codes are from CoSQA+, while 39.7% are from CoSQA.
 This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes.
 As shown in Table III, the three models fine-tuned with CoSQA+ outperform those fine-tuned with CoSQA, which indicates better training performance of CoSQA+.
-Notably, the improvement in NDCG@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
+Notably, the improvement in MAP@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
 RQ1 Summary: CoSQA+ has better quality than CoSQA. Fine-tuning CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ yields better performance than with CoSQA, especially for multiple relevant codes.
@@
 Experimental Setup
-To establish a reliable ground truth, we first have three Python experts design test programs for 400 randomly selected query-code pairs and annotate them based on test execution results.
-The final ground truth labels are determined by majority voting among these three experts.
-We then evaluate three different annotation approaches against this ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet).
-For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators [61].
+To establish a reliable ground truth, we recruit three experienced Python developers with 3-5 years of professional programming experience and computer science backgrounds to serve as expert annotators.
+These experts first independently design test programs for 1000 randomly selected query-code pairs from CoSQA+, following a standardized evaluation protocol that required: (1) Functional correctness verification through executable test cases.
+(2) Edge case coverage assessment.
+(3) Query requirement satisfaction analysis.
+They then annotate the selected pairs based on test execution results.
+The final ground truth labels are determined by majority voting among these three experts, with an initial inter-annotator agreement measured by Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators [63].
+In cases of disagreement, experts engage in discussion sessions to reach a consensus.
+We then evaluate three different annotation approaches against this established ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet), each provided with identical query-code pairs and evaluation criteria.
 Results and Analysis
@@
 Method Accuracy
-Agent-based Annotator 0.920
-Human Wo. Test Program 0.880
-Deepseek-V3 0.865
-O1-mini 0.833
-Claude-3.5-Sonnet 0.845
-As shown in Table IV, our agent-based annotator achieves 92.0% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (84.0%), O1-mini (82.0%) and Claude-3.5-Sonnet (84.0%).
+Agent-based Annotator 0.939±0.020
+Human Wo. Test Program 0.891±0.026
+Deepseek-V3 0.883±0.027
+O1-mini 0.881±0.027
+Claude-3.5-Sonnet 0.863±0.029
+As shown in Table IV, our agent-based annotator achieves 93.9% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.1%), Deepseek-V3 (88.3%), O1-mini (88.1%) and Claude-3.5-Sonnet (86.3%).
 This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification.
+To ensure appropriate interval estimation for our data characteristics, we adopt the Clopper–Pearson exact method with 95% confidence intervals for accuracy metrics in TABLE IV.
+Clopper–Pearson exact method is well-suited for moderate sample sizes and high sample proportions (i.e., proportions close to 0 or 1).
 Importantly, about 83% of the test programs generated by our agent are successfully executable, indicating the reliability of our test program generation approach.
@@
 The superior performance of our agent-based annotators can be attributed to two key factors: 1) The test program provides concrete evidence of whether the code satisfies the query requirements, reducing ambiguity in the annotation process; 2) The systematic verification through test execution helps eliminate human bias and inconsistency that often occur in manual annotation.
-In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.818, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
-RQ2 Summary: Our agent-based annotator achieves 92.0% accuracy, outperforming both traditional human annotation and single LLM approaches.
+In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.82, notably higher than human annotations without test program verification (0.60), which suggests that test programs help reduce subjectivity in the annotation process.
+RQ2 Summary: Our agent-based annotator achieves 93.9% accuracy, outperforming both traditional human annotation and single LLM approaches.
 The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
@@
 Experimental Setup
-The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using NDCG@10 as the primary metric.
-NDCG@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match.
+The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using MAP@10 as the primary metric.
+MAP@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match.
 Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly.
-To further comprehensively evaluate the performance of the models, we also introduce Mean Average Precision (MAP) and Recall as additional metrics.
-MAP provides a measure of the overall ranking quality by averaging the precision scores across all relevant results, which is especially useful for scenarios where the order of multiple correct matches matters.
-Meanwhile, Recall evaluates the model’s ability to retrieve all relevant code snippets from the dataset, ensuring that no important results are missed.
-Together, these metrics offer a more robust assessment of the model’s effectiveness in semantic code search tasks.
+To provide a more comprehensive evaluation, we also include NDCG [64, 22, 65, 66] (Normalized Discounted Cumulative Gain) and Recall as additional metrics.
+Although NDCG is typically used for graded relevance judgments, in our case where relevance is binary (either matching or not), it still serves as a useful metric by emphasizing the importance of ranking relevant results higher in the list.
+Meanwhile, Recall measures the model’s ability to retrieve all relevant code snippets from the dataset, ensuring comprehensive coverage of correct results.
+Together, these metrics offer a robust and multi-faceted assessment of the model’s performance in semantic code search tasks.
 The tested methods include both traditional information retrieval techniques and advanced deep learning-based models.
@@
 This represents a significant advancement in embedding models in recent years, providing a solid foundation for the practical implementation of semantic search based on large models.
-From the evaluation results, the performance differences among embedding models seem not significant.
-However, due to the richness of CoSQA+, which effectively eliminates random factors, even subtle differences in metrics can reflect the overall capability gap between models.
+Regarding the evaluation metrics (MRR, NDCG, MAP): these inherently assign high weight to the top-ranked results.
+Consequently, the most critical factor influencing a model’s score is its ability to rank correct candidate snippets at the very top positions.
+The internal relative order of matched code snippets in the search results has no influence on the score.
+A model is considered better on CoSQA+ if it can find more matched code and position them at the top of the results.
+From the evaluation results, the performance differences among embedding models do not seem significant.
+However, due to the richness of the CoSQA+, which effectively eliminates random factors, even subtle differences in metrics can reflect the overall capability gap between models.
 Moreover, embedding models generally perform poorly on our benchmark.
@@
 RQ3 Summary: The evaluation on CoSQA+ clearly indicates that deep learning-based models significantly outperform traditional retrieval methods for semantic code search.
-Notably, CodeBERT leads the performance ranking by achieving the highest scores in both NDCG@10 and MRR, while the alternative embedding-based models also offer competitive results.
+Notably, CodeBERT leads the performance ranking by achieving the highest scores in both MAP@10 and MRR, while the alternative embedding-based models also offer competitive results.
 V-D RQ4: Contributions of Different Components in Test-Driven Agents
@@
 Configuration Accuracy
-Single LLM 0.865
-PS + FA 0.840
-PS + TPG + FA 0.833
-PS + TPG + TE + FA 0.835
-PS + TPG + Simulated Executor + FA 0.840
-PS + TPG + TE + BF + FA 0.920
+Single LLM 0.883
+PS + FA 0.873
+PS + TPG + FA 0.852
+PS + TPG + TE + FA 0.854
+PS + TPG + Simulated Executor + FA 0.853
+PS + TPG + TE + BF + FA 0.939
 TABLE VI: Accuracy of Various Test-Driven Agent Configurations (Abbreviated)
-The results indicate that the baseline Single LLM achieves an accuracy of 0.865.
-When the system uses only the Preliminary Screener in combination with the Final Arbiter, the accuracy slightly drops to 0.840.
-Incorporating the Test Program Generator alongside these yields an accuracy of 0.833, suggesting that generating test programs without execution may introduce noise or irrelevant information, which can harm performance [62].
+The results indicate that the baseline Single LLM achieves an accuracy of 0.883.
+When the system uses only the Preliminary Screener in combination with the Final Arbiter, the accuracy slightly drops to 0.873.
+Incorporating the Test Program Generator alongside these yields an accuracy of 0.852, suggesting that generating test programs without execution may introduce noise or irrelevant information, which can harm performance [67].
 Adding the Test Executor to the configuration improves the performance marginally (0.854), and using a Simulated Test Executor does not provide a significant advantage (0.853).
 This aligns with the observation that simulated execution cannot fully capture the complexity of the real compile-and-run process, potentially introducing inaccuracies or misleading feedback.
-Notably, the inclusion of the Bug Fixer increases the accuracy substantially to 0.920.
+Notably, the inclusion of the Bug Fixer increases the accuracy substantially to 0.939.
 This enhancement demonstrates that the Bug Fixer is crucial for processing the test programs effectively, as it enables the system to generate real execution results that are vital for improving prediction accuracy.
@@
 Simulated execution results cannot replace the benefits of real execution feedback.
+V-E RQ5: Cross-Language Generalization of the Test-Driven Agent
+This experiment is to explore the generalizability of the test-driven agent’s annotation performance beyond Python.
+Experimental Setup
+We collect and annotate query-code pairs from CodeSearchNet (CSN) for PHP, Java, and Go to benchmark the test-driven and single-LLM annotation methods.
+We select these languages because (1) they rank among the top ten most-used languages, ensuring real-world relevance [68], and (2) they represent distinct paradigms: dynamically typed scripting (PHP), statically typed object-oriented (Java), and statically typed procedural/concurrent (Go).
+We build the benchmark in two steps.
+First, we extract unannotated query-code pairs from CSN for each language.
+Second, we label every pair using the same pipeline described in RQ2 to ensure consistent evaluation.
+To evaluate cross-language performance, we adapt the pipeline in Algorithm 1 to each language.
+We slightly refine the prompts and configure separate execution environments for PHP, Java, and Go.
+We then compare two annotation methods against the human-labeled ground truth: (1) the test-driven agent based on DeepSeek-V3 and (2) the single-pass LLM annotator (DeepSeek-V3).
+Accuracy is the metric, defined as the fraction of labels that match human annotations.
+Result and Analysis
+The results reveal a consistent pattern: the test-driven agent achieves higher annotation accuracy than single LLM annotators across all three non-Python languages, with significant improvements observed in PHP (82.89% vs.
+72.70%, +12.17%), Java (79.96% vs.
+69.59%, +10.37%), and Go (81.93% vs.
+73.49%, +8.44%).
+This performance improvement indicates that the test-driven framework, by incorporating functional validation and error correction mechanisms, effectively enhances cross-language code understanding capabilities, outperforming the baseline of single LLMs.
+Our test-driven agent consistently outperforms the single LLM baseline in accuracy.
+This highlights a key advantage: our approach mitigates the limitations of pre-trained models like DeepSeek-V3.
+Such models may exhibit inherent limitations stemming from their training data composition, which often over-represents certain languages or paradigms.
+This can reduce their robustness when handling syntax and features specific to other languages.
+For instance, PHP’s dynamic typing quirks pose interpretive challenges for the single LLM baseline.
+However, the test-driven agent’s structured validation process effectively addresses the issues, resulting in clearer improvements.
+Notably, PHP shows the largest accuracy gain (12.17%) compared to the baseline, which could stem from its unique blend of scripting and web-focused features—elements that demand stricter contextual validation.
+RQ5 Summary: The test-driven agent outperforms single LLM annotators across PHP, Java, and Go with an average 10.33% accuracy gain, showing strong cross-language generalization.
 VI Discussion
@@
 This conclusion aligns with intuition, but with the analysis of other datasets and the construction of CoSQA+, we find that the decline in performance is not only due to the need to exclude more confusing codes but also because of the increased likelihood of test misjudgments.
-For example, a LLM may select a piece of code that meets the query requirements, but this query-code pair is not included in the candidate code set.
+For example, an LLM may select a piece of code that meets the query requirements, but this query-code pair is not included in the candidate code set.
 As a result, this correct selection is misjudged as incorrect.
@@
 We encourage the community to adopt ethical guidelines when using our benchmark and tools, ensuring that their application aligns with positive social impact.
+VI-D Failure Case Analysis
+Our test-driven agents achieve 93.9% accuracy, but understanding the remaining 6.1% failure cases is crucial for future improvements.
+We analyze ten false-positive and false-negative cases and identify four primary root causes: (1) Ambiguous queries (30% of cases) where natural language descriptions like "python fabric verbose logging" have multiple valid interpretations (e.g., generic logging functionality vs.
+Fabric library-specific SSH logging), leading to mismatched test generation; (2) Human annotation biases (30% of cases) where annotator subjectivity leads to incorrect judgments, such as being misled by superficial keyword matching while agents correctly identify functional mismatches; (3) Semantic misalignment (20% of cases), where agents fail to capture the precise intent behind technical terms or behavioral descriptors; (4) Validation misalignment (20% of cases), where test programs validate code functionality rather than query requirements, such as testing all features of a function instead of the specific query need.
+The analysis reveals that most failure cases stem from inherent limitations in natural language understanding rather than systematic flaws in our test-driven approach.
+Query ambiguity remains the most significant challenge, suggesting that future work should focus on multi-interpretation test generation or query clarification mechanisms.
+Importantly, our agents sometimes perform better than ground truth labels, indicating that test-driven verification can identify annotation failure cases in human-labeled datasets.
+VI-E Validation Strategy and Scalability Constraints
+Complete human verification of large-scale datasets faces fundamental scalability barriers.
+Expert annotation with test programs requires 34 minutes per pair, making exhaustive validation of 412,080 pairs economically prohibitive at approximately 234,000 person-hours.
+Our sampling-based validation strategy addresses this constraint while maintaining methodological rigor.
+To strengthen validation beyond our 1,000-sample ground truth evaluation, we implemented an additional verification mechanism: a specialized checker agent using DeepSeek-V3 that performs secondary validation of our test-driven annotations.
+This checker evaluates the complete annotation pipeline output, including preliminary screening results, test programs, execution outcomes, and final judgments.
+The checker achieves high agreement with human expert annotations (Krippendorff’s $\alpha$ = 0.914) and, as ground truth, grants the test-driven agents 92.8% accuracy.
+The high inter-annotator agreement between our checker and human experts demonstrates that automated verification can reliably assess annotation quality at scale.
+This checker-based validation mechanism enables quality assurance across the entire dataset while maintaining economic feasibility, offering a practical solution to the scalability constraints of traditional manual verification approaches.
 VII Conclusion
@@
 Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA.
-The test-driven annotation approach achieves a high accuracy of 92.0%, outperforming traditional human annotation and single LLM methods.
+The test-driven annotation approach achieves a high accuracy of 93.9%, outperforming not only traditional human annotation and single-LLM methods but also demonstrating strong cross-language generalizability.
 Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks.